A Bibliography of Publications in Byte Magazine: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 January 2024 Version 2.15

Title word cross-reference

1-2-3 [Ren89]. 1.0 [Woo88b]. 1100/33 [Hol89d]. 16-bit [Kli89a]. 1986 [Wit87b, Wit87a]. 1988 [Vos89].

2 [AD88c, Gil88, JM84, Kli89c, MS88, Min88, Min89a, Min89c, Min89b, Min89d, Min89e, Nan89, Sch88, Sch89d, Thu89, Wir84]. 2.0 [Sha89c]. 220 [Hol88c]. 286 [Van89a]. 286L [Ung89c]. 2s [Smi89a].

3-D [Kli89b, LG88]. 310 [Ung88a]. 32-Bit [Tho89c]. 386 [Hol88b]. 386/32 [Api89]. 386s [Hol88d]. 3D [HU89].

80 [CC80, Cra89c, GR81]. 80386-based [AD88a]. 80386s [AD88a, Mal89]. 8514/A [Kli89c]. 86 [Sch86a].

= [Rob89d].

ABCs [Mee89a]. ability [Fid88]. Abstract [Per86, Way88a]. access [Che89, Fid88]. accidental [AK89]. Acer [Hol89d]. Action [Lan89c, Win88]. Ada [NU89, Sha88a]. Adapter [Wil85]. ADC [Sat89a]. Adding [Bat88]. Addison [Wit87a]. Addison-Wesley [Wit87a]. Advanced [Hol89a, Vos89]. after [Sha88c]. age [AD88d, Cra89e]. Agenda [Woo88b]. agents [CM88]. AI [Dee85, Rob87b]. aid [Tho89g]. AIX [Smi89a]. Algorithm [Phi86, RS87a, BRS80a, BRS80b, RS87b]. Algorithms [Jos87]. Alpha [Rub89b]. ALR [Van88a]. alternative [CV89]. Alto [Wad81]. always [Paj89]. Amazing [Wit87a]. American [Van89a]. Analysis [Mac89, Ngu81]. Anatomy [VC89a]. ANSI [Vos89]. any [Dro89]. APPC [Dav89a]. appeal [Hol88d]. Apple [Bar89a, SS80, Tho89c, Wil84]. application [Sch86b]. applications [Gla89d, Sch89b]. Approach [Lan88b]. Approaches [AR85]. Approximations [Mos86]. APT [Way88a]. Arabic [Mac86]. Architectures [Dee85]. aren’t [Paj89]. array [Bat88]. arrives [Bar89b, Rub89a]. Art [Per86, Rob89d]. Arthur [Ano86a]. Artist [Kli89c]. Arts [Wit87b]. assembler [Sat89b]. assembler/disassembler [Sat89b]. assemblers [Bla89]. assembly [Gre88c, Gre89g]. assistance [Gre89b, Gre89c]. Associative [Kos87]. AST [Hol88a]. AT-Compatibles [Wil88c]. ATOMCC [Cha86]. Attitude [Lin88]. Automatically [JS85]. Autonomous [JHW86]. avoids [JM84].

Back [JH87, Cra88]. Back-Propagation [JH87]. Backup [Wal88b]. backups [Cra89a]. bandits [AK89]. Banzhaf [Sch84]. barrier [Hel89a]. base [Lan89b]. based [AD88a, Die89]. BASIC [Bas84, Sha88b]. Battle [ADG89a, Hay89a]. BDOS [Pou82]. Be [Rob87a, Cra89h, Obe88]. Beefed [Mee88a]. Beefed-up [Mee88a]. beefs [GT88]. Behind [Ing81, AEG89]. Benchmark [Kar85]. Benchmarks [AD88b]. best [Paj89]. Better [Str88, BS88]. Between [Tel88, Bry89a]. Beyond [Kur89, Mia89a]. Big [Tho89d, Nic88]. bigger [Ras89b]. BIOS [Mye88a, Gla89c]. Bit [Fre86, Han87, Tho89c, Kli89a]. Bit-Mapped [Fre86, Han87]. bits [AD89a]. block [Wal85]. blocks [WH89]. Blues [Pou82, ADG89b]. board [Hol89b, Kli89d, Mal88b, Mee88b, Pou89c, WN88]. Boards [Wil88c, Kli89c, Mee88a, Ras89d, Tho89f]. booster [Tho89g]. Borland [GT88]. box [Tho89d]. Brain [Ste85, Zei87, Fid88]. brains [AEG89]. brave [Pas89]. Breaking [Hel89a]. bright [Ras88b]. Bringing [Loe88]. brownout [ADG89b]. Bucking [Ung88a]. Build [Cia86]. Building [WH89, WH87, Wal85]. bulletin [Mee88a]. bus [AP89, New88, Whi89]. Business [Mac89]. buyer [Ano89b, AD88d, DA88, Mal88b, Mee88b]. BYSO [Won86]. Bytelines [Lib85].

C [Vos89, AU89, Bat88, Edg87, Hom89, KR88, Lin88, Sha88e, Str88, Ude88]. C-talk [Sha88e]. cables [Bry89a]. cache [Hol89b, Van88a]. Caching [Gla89a]. CAD [Tut89]. cafeteria [Van89a]. can [Cra89h, Mee88d, Obe88]. can’t [Mee88d]. Carlo [Mac89]. Carnegie [Han87]. Carnegie-Mellon [Han87]. CAS [Gla89d]. case [Ano89b, Ano89b, Gib89, Jon89, McC89, Sch89c]. CD [Byt88c, Ore88]. CD-ROM [Ore88]. CD-ROMs [Byt88c]. cellar [Cia88a, Cia88b]. Challenge [Hew85]. Changed [Tou88]. Changing [Daw89a]. channel [Van89c, New88]. cheap [Cra89h]. checking [Lan89b]. Cheetah [Nad89]. Children [Wit87a]. chip [Hay89b]. Chips [Mar84, Hay89a]. choice [Dav89a]. choose [Bry89a, Jon89]. Ciarcia [Cia88a]. circuit [Cia88a, Cia88b]. Claris [Tut89]. Clash [Coo89, Ude89a]. Classifier [Fre86]. Classroom [Wit87b]. clean [Gre89f]. climbing [Api89]. Clipper [Sch89b]. clone [Hay89c]. clones [AD88a]. Code [Rob87a]. Codes [Dro89]. Coding [Ams86]. Color [Qui89a, Egl89, Gro80]. colors [Tho89f]. columns [Lic89]. combined [Hol88b]. come [AD88d]. Comes [Wil88c, Sha88a, Sha89a, Val89c]. communications [CS88, Gil88, Mee89b]. communicator [Bar89c]. compact [Bar89a]. companion [Loe89a]. Compaq [Hay88a, Hol88d, Hol89c]. Compatibles [Wil88c, Won86, Bro87]. compiler [Sha88a]. Compilers [Rob87a, AU89]. compress [BS88]. Compression [Ams86]. CompuStar [Van89a]. Computer [BD84, Bro84, Ken89, Ker85, Mac86, Mos86, Mye88a, Vos89, Wad81, Wil83, Wil84, Wit87b, Won86, BL80, Cla89, Cra88, Gre89a, Koc89, LG88, Loe89a, Obe88, Pou89c, TB88, Tho89h, Wil88a, Knu86]. computerisation [Gig89]. Computers [Hoe87, Pou85a, Wit87a, Ras88c, Ras89a]. Computing [Ras89a, Van89a]. Configuring [Pou89e]. Connection [Hay88b, Ore88, Smi89c, Tit89]. Connections [Fel85]. considerations [Knu86]. Constraint [Las87]. Constructing [Kos87]. Consultants [Pou85a]. control [CS88, MT89a, Qui89b]. controllers [Hol89a]. converge [Bar89h]. coprocessing [Fri89]. coprocessor [Gre88a, Gre88c, Kli89c, Kli89d, Pou89c, Rob87b, Tho87]. Corel [Ros89]. Corp [Mye88a]. Could [Rob87a]. counting [Ste88]. couple [Ras88c]. covers [Tho89c]. CPU [Pou88b, Pou88e, Van88b]. Cray [Hay89b]. Cray-on-a-chip [Hay89b]. Creating [MT87]. Critchlow [Ano86a]. crop [Nic89a]. Crossbar [Hay88b]. Cryptographic [Smi83]. Cryptography [Smi83]. Culture [Ker85]. Curing [ADG89b]. current [Nic89a]. Cycle [WH87].

D [Kli89b, LG88, Opp88a]. damned [Pea88]. dashboard [Min88]. Data [Ams86, Cia86, Han87, LM85, Opp88a, TT86, Woo88a, AK89, Ano89a, Dro89, Ngu81, vNC89]. Database [Woo88b, Rob89b, VC89f, vNC89]. databases [Dav89b, Daw89b]. DataPlex [Woo88a]. Dayflo [Sha84]. dBASE [Rub89a, VC89f]. DBMS [Ano89a, MB89, Opp88a, Pas89, Rub89b, Bar89f]. Dead [Ung89a]. Dealing [MHKO89]. debugger [Sha88f]. debugging [Sha88f]. Debugs [JS85]. Debunking [Kli89a]. DEC [SM89]. Declarative [ES85]. delivery [Sch89a]. Dell [Hol88c, Ung88a]. delta [JH87]. description [JM84]. Design [AS88, Ing81, JM84]. Designing [Duf87, SIK^{+}82]. desktop [Sat89a, Ste89b]. destruction [AK89]. details [DA89]. development [PK88, Tho89h]. device [WA89]. Dewdney [Ken89]. diagnostic [Gre88d]. Dialing [Mee89b]. dictionary [Gig89]. different [Sch89c]. Digital [Gla89h, Pou89a, MHKO89]. digitizing [DA89]. dilemma [Hol89c]. dimension [HU89, LG88]. dimensional [SS80]. dimensions [Bat88]. Directory [Gre89b, Gre89c]. disassembler [Sat89b]. disc [Gla89g, Gre88d]. disk [Gla89b, Hol89a, Min89b]. display [PKAG88, SB88, Smi89b]. Displays [Han87, SD88]. distributed [Dav89b, Hou89, MT89a, WSW89, WP89]. Do [Ras88a, Ras89e]. document [Sch89a]. Dolch [Van89b]. Domesticating [Lan89a]. done [Bro84, Gel88]. Downsizing [Sat89a]. Draw [Ros89]. drawing [DeM88]. drive [Cra89a, Hol89a]. drivers [WA89]. DSI [PK88]. DSPs [Saf89]. DTP [Gab89b]. duelling [Lan89b]. Duncan [Bou88].

E-mail [Mee89c]. Ease [Val89a]. eases [Cra89a]. Easy [Rob89a]. economical [Nad89]. economics [Mee89c]. editing [Daw89a, Sha88d]. Edition [Vos89]. Editor [Han87, Loe89b]. editors [Ude89b]. Education [Wit87a]. Effectiveness [AR85]. Efficient [Dee85, Duf87]. Efforts [JS85]. EISA [Bar89b]. Electing [Min89a]. Electronic [Coo84, Gig89]. Embedded [Bet88]. EMS [Hol89b]. encryption [Dro89]. Encryptor [Cia86]. Encyclopedia [Bou88]. Encyclopedias [Coo84]. end [Gla89i, Lan89b]. enemy [Gre89a]. Engineering [Ste85, OGY^{+}89]. English [Wit87b, Byt88c]. English/Language [Wit87b]. Enhanced [Wil85]. enigma [MB89]. enough [Sta89]. entry [Woo88a]. Environment [Tes81, Gre81, Hol88b, NU89, Sha88d, WP89]. Equations [Ste87]. Error [Way88b]. Error-Free [Way88b]. errors [AD89a]. Ethernet [Bir83, Mie84]. evaluated [Mee89d]. Events [Sch86c]. Ever [BN89, Mal89]. Ever-expanding [BN89]. Evolution [Mie84]. Excel [Kyd88]. Excursions [Ken89]. expanded [Yan89]. expanding [BN89]. Experience [Jos87, McC89, Ung89b]. Expert [Fic85]. experts [OGY^{+}89]. Exploring [Sch88]. Extend [Edg87, Val89b]. Extensible [Ude89b].

Face [Hoe88, SB88]. faire [Mee88b]. family [Daw89b]. fashion [Smi89b]. Fast [O’N88, Rob87a, Tho87, Ano89c, Kin89]. Faster [O’N88, Sna87]. fastest [Mal89]. Fax [Mee88b, Cra89d, Ras89d]. Features [Gab88, Che89]. few [Ras89d]. Fiber [Bry89a]. fiber-optic [Bry89a]. file [Ano89a, Che89, Mee89d, Eug89]. filenames [Ker89]. Finding [TT86]. fired [OB89]. First [Sha84, Tho87]. Flaws [JM84]. Floating [Gre88a, Gre89d, Kar85, Wil88b, Gre88c]. Floating-Point [Gre89d, Kar85, Wil88b, Gre88a, Gre88c]. floppy [Hol89a]. flying [Cra89b]. focus [DA89]. forces [Sch89d]. forever [HNM89]. formats [Gra89]. Four [Rub89b]. Fourier [O’N88]. FP [HK85]. Fractals [MT87]. Fractions [Way88b]. frame [WN88]. framework [Sch86b]. Free [Way88b]. Full [Tho89e, Gab89a]. Full-spectrum [Tho89e]. FullWrite [Gab88]. Functional [HK85]. Future [FP89, BL80, TS89].

G. [Knu86]. gateways [Mee88d]. Gehani [Vos89]. general [CC80]. generalized [JH87]. Generating [Whi84]. Generator [WH87]. Generators [MSC87]. geography [Min89b]. get [Mee88d, OB89, Sch89b]. Getting [Gel88, Min89c, Ras89b]. glance [AD88b]. Goals [Wir84]. goes [Cra89d, Woo88a]. Gold [Nad89]. good [Dro89]. grabber [WN88]. Grand [Fid88]. Graphic [Gro80, DA89, Rob89b]. graphic/database [Rob89b]. graphical [HB89a]. Graphics [Gra89, Nic88, Wil85, AEG89, Bar89f, BL80, CC80, Coo89, Kli89b, Kli89c, Kli89d, MS88, Ros89, SS80, Tho89f, Kli89d]. Greased [Ude89c]. great [Bar89c, Ros89]. Gregg [Knu86]. Groping [Ras89c]. group [Ras89c]. groupware [CM88, Gru88, Opp88b, Ras89c, Win88]. groupwise [EL88]. Growing [Bry89b]. grows [Mee88c]. Guide [Ano88, Her87, Wit87b, Ano89b, Ano89e, AD88d, DA88, HB89a, Mal88b, Mee88b]. GUIs [HB89a].

Hall [Wit87b]. hand [Ras89a, VC89b]. hand-held [Ras89a, VC89b]. Hands [FHMC89]. Hands-on [FHMC89]. Handy [VC89b]. Hard [Gla89b, Gla89g, Gre88d]. Hardware [Cia86, Dee85, Sar89, Sha88f]. hardware-style [Sha88f]. hashing [Sna87]. healthy [Wil88a]. heat [Ung89a]. heir [Way89a]. heirs [Mee89d]. held [Ras89a, VC89b]. helical [Bre89]. helical-scan [Bre89]. help [Ras88a]. here [Mee88d]. here-or [Mee88d]. Heuristic [Jos87, Ste87]. Heuristics [Jos87]. HFs [Gre89c]. hidden [TP89]. High [Smi89b, Van89a, Woo88a]. High-fashion [Smi89b]. High-tech [Van89a, Woo88a]. History [Wir84]. hits [Thu89]. HM [Pou89b]. Hooked [Cra89c]. HOOPS [Kli89b]. Hope [Bai85]. HP [Bar89b, Smi89b]. Huffman [Ams86]. Human [AR85]. hybrid [Lan89c]. Hyperquotes [Byt88a]. Hypertext [Byt88a, Byt88b, Byt88d, Fid88, BC88, Fri88].

IBM [Mye88a, Mye88b, Bir83, BD84, Bro87, Die89, Gla89c, Hal89a, Han87, Kli89c, Lev88, MB89, Mye88a, Tit89, Wei84, Won86]. ICE [Sha88f]. Idea [Her85]. If [Gre89e]. II [BRS80b, DeM88, Pou89f, SS80]. III [Col87, Gre89h]. IIx [Bar89a]. image [Ano89e, Daw89a]. images [BS88, SP89]. Immense [Nel88, Mun89]. Impact [Ker85, Gab89a]. imperfect [FP89]. Implementations [Won86]. Improved [Rob87a]. improves [Van88a]. Index [Sch84]. Information [Bha87, Fid88]. Input [Gre89f]. Installing [Edg87]. IntegrAda [NU89]. Intel [Hay89b, Mar89]. Intellectual [Ker85]. Intelligent [CM88]. Interacting [Mun89]. Interactive [Bor87, Gla89h]. Interface [Mye88a, SIK^{+}82, Apo89, Cra89e, Hoe88]. interfaces [DHM89b, Gla89b, HB89a, Tel88]. Interleaf [Ude89d]. International [Sch86c]. interpolating [CC80]. Interpretation [PW86]. interpreter [Sch88]. Interrupts [Paj89]. interview [Knu86]. Intrinsic [MSC87]. Introduction [BW88, Mac89, Smi83, Vos89, Ano86a]. intuitively [Fid88]. Inversion [Phi86]. issue [Ano81, Ano86b]. IV [Rub89a].

J [Ano86a]. job [BC88, Gel88]. join [Sch89d]. Just [Ras89d].

Karmarkar [RS87a, RS87b]. Keep [Wil88a, Gre89f]. Keeper [Wei84]. Keeping [Van88b]. Key [Smi83]. keys [Gre89g, Gre89i, Gre89h]. Khachiyan [BRS80a, BRS80b]. kind [Sch89c]. king [Hal89a]. Kit [Wil88b]. Know [Gre89a]. Knowledge [Way87]. Knuth [Hoe87].

Lab [Bar89i]. LAN [AD88c, ADG89a, CV89, Gla89i, Ras89e, Saa89, Sta89, Tho89g, VC89a, VC89c, Wat89]. Language [AS88, Duf87, Kae81, Obe87, Opp88a, PW86, Wit87b, Bas84, Gel88, Pou84b, Qui89b, VC89f, Weg89]. Languages [ES85, Bet88, GT88, Qui89b, Weg89]. LANs [Ras89b]. Laptop [Hol89c, Hay88a, Loe89a, Ras88b]. Laptops [HNM89, BN89, Ung89c]. Large [Phi86, AD89b, RS87b]. large-screen [AD89b]. lasers [Qui89b]. last [Cra89g, Hay88a]. layers [TP89]. Learn [Jos87]. Learning [Hin85, Hin86, Weg89, JH87]. letter [Phi87]. Lexical [Won87]. Lies [Pea88]. Life [Sha88c]. Light [Ras88b, Gla89i]. Lighting [SD88]. lightning [Ude89c]. limit [Eug89, Hay89d]. Linda [Gel88]. linear [BRS80a, RS87b]. lineup [Hal89b]. links [Cra88]. Lint [Lan88a]. Lisa [Wil83]. lis’ner [Cia84]. LISP [AS88, AR85, Ano88, BW88, BD84, JS85, LB81, Tou88, Win85, Won85, Won86, Won87, Lan89c, Sch88]. LISPs [Pou82]. Logic [Lan88b, Las87]. Logical [Wei84, Dav89a]. Long [WH87, Hay88a]. Look [Sha84, Sna87, JM84, Tho87, WN88, Hoe88]. Looking [Mia89a, Saa89]. Lotus [Ren89]. LU [Dav89a].

Mac [Cra89d, Ste89b, Cra89c, Bar89a, Bar89d, Cra88, Cra89a, Cra89e, Sha88a, Tho89d, Tho89g, Val89a, Val89c, Val89d, Way89b]. MacApp [Sch86b]. MacDraw [DeM88]. Mach [TS89]. machine [Mal87, Tel88]. Macintosh [Ben85, Cra89f, Gre89c, Lan89c, Loe88, Loe89a, Loe89b, Sha89a, Sha89c, WA89, Wil84]. Macintosh/Lisp [Lan89c]. Macintoshes [Wil88c, Tit89]. madness [Ano89c, Sha89c]. mail [Mee89c]. maintenance [Gla89g, Gre88d, Wil88a]. Making [Ano89b, Gla89d, Tit89, Wll89]. Man [Ung89b, Tel88]. Manage [Woo89a]. management [Gre89e, Woo88c, Woo89a].

Manager [AD88c]. Managing [Nel88]. Mandelbrot [Pou86b, Sch86d]. Mankind [Byt88d]. many [AP89]. Mapped [Fre86, Han87]. march [Kli89c, Yag89]. masses [AD88a]. masters [AP89]. match [Bry89a]. Math [Tho87, Way89b]. Mathematica [Way89b]. Mathematical [Per86, Ste87]. Mathematics [Hoa86]. Matrices [Phi86]. maybe [Ras89e]. MC [Kli89c, Van89c]. MCA [VC89d]. Mechanic [Fic85]. Media [Rot83, Bar89i]. Meets [Bir83, DO88, VC89d]. Megahertz [Ano89c]. Mellon [Han87]. Memory [Edg87, Kae81, Kos87, Mal88b, Mye88b, Rot83, Gla89a, Gre89e, Hel89a, Hol89b, Yan89]. Memory-Resident [Edg87]. Merge [Hol88b]. metal [Bry89a]. metal-mix [Bry89a]. metallic [Bry89a]. method [RS87b]. Methodology [OGY^{+}89]. Methods [Ste87, Gla89j, Ngu81]. MHz [Api89]. Michael [Knu86]. micro [Hol89d, New88]. Microcomputer [LB81, Ste89a, Ung89a]. Microcomputers [Pou84a, Smi83, Ano89d, AD88b, Hol88d]. microcontrollers [Cia88a]. MicroExplorer [Lan89c]. Microsoft [Lan89a]. million [Tho89f]. mimics [Fid88]. Mind [Ben85]. Minicomputer [Pou89c]. Minstrel [Pou89b]. Mitsubishi [Ung89c]. mix [Bry89a]. mod [Hel89a]. mode [Gla89k]. model [TS89]. Modeling [Zei87]. modelling [HU89]. models [Daw89b]. Modem [HS89, Gla89j]. modems [AD89a]. Modern [Gla89j]. modes [Wil88d]. Modula [JM84, JM84, Nan89, Sch89d, Wir84]. Modula-2 [JM84, JM84, Nan89, Sch89d, Wir84]. modules [Tho89g]. molecular [Cla89]. monitors [AD89b]. Monte [Mac89]. Most [Wit87a]. Motorola [Tho87]. Move [Gab89b]. MP [Ung89c]. MP-286L [Ung89c]. MS [Bou88, Bla89, Gre89c]. MS-DOS [Bou88, Bla89, Gre89c]. MuLISP [Sch86a]. MuLISP-86 [Sch86a]. Multiple [Kyd88]. multitasking [Min88, Min89d, Sha88c]. must [Dro89].

’n [Gre89g, Gre89h]. Natural

[Obe87, PW86]. Navigation [JHW86]. NEC [VC89e]. need [Ras89e]. neither [MB89]. Nelson [Byt88a]. NetBIOS [Gla89f]. Network [Jos87, ADG89a, Bry89b]. Networking [Com89a]. Networks [Hin85, Hin86, OB89, Thu89, WH89]. Neural [Jos87, OB89, WH89]. Neural-Network [Jos87]. Next [Tho89b, Tho89h, TB88, Tho89h]. NextStep [Tho89h]. nimble [Hol88c]. no [AD89a, Rub89b]. nor [MB89]. Novice [JS85]. Number [MSC87, WH87]. Numbers [Whi84, Qui89a]. numeric [Fri89]. Numerical [Ngu81].

Object [Ano86b, Kae81, Pou88e, Rob81, Daw89b, DHM89b, Lin88, Pou88b, TS89, Tho89a, Tho89h, Ude89a, Weg89]. Object-Oriented [Ano86b, Kae81, Pou88e, Rob81, Daw89b, DHM89b, Lin88, Pou88b, TS89, Tho89a, Tho89h, Ude89a, Weg89]. Occam [Pou84b, Pou89f]. odd [Ras88c]. off [Bar89d, PKAG88]. office [Hou89]. Ogivar [Hol89c]. old [Bre89]. Omnibus [Ken89]. One [Ung89b, Sta89]. only [Obe88, Ren89, Van89c]. Open [Hew85, Hoe88]. Opening [Col87]. operating [ADG89a, Gre81, Sha88c, TS89, VC89a]. optic [Bry89a]. Optical [Rot83, AE89]. Optical-Memory [Rot83]. Optimizing [Fri89, Rob87a, AU89]. option [AE89]. Opus [Rob89b]. Oriented [Ano86b, Kae81, Pou88e, Rob81, Bar89f, Daw89b, DHM89b, Lin88, Pou88b, TS89, Tho89a, Tho89h, Ude89a, Weg89]. OS/2 [AD88c, Gil88, MS88, Min88, Min89a, Min89c, Min89b, Min89d, Min89e, Sch88, Sch89d, Thu89]. OSI [VC89c]. Overfitting [LM85]. overlays [Gre88b]. Overview [ES85, Gre88b]. Oxford [Gig89].

PAC [Van89b]. pack [Mee89d]. package [CC80, Gab89a, Gab89b, Lan89a, Sha88a, Tut89, Woo88a, Woo89a]. packages [HU89, Mia89b]. packet [Sat89b, Wll89]. pads [Sat89b]. page [Rob89a]. page-recognition [Rob89a]. PageMaker [Gab89b]. painlessly [VC89e]. pains [Bry89b]. pair [Ung89c]. Palmer [Wit87a]. Pan [Phi86, Phi87]. Pan-Reif [Phi87]. paper [Pou89a]. paperless [Hou89]. paperwork [Sch89a]. Paradox [MB89]. Parallel [Hin85, Hin86, Pou85b, Pou89e, Wil88c, FHMC89, Gel88, Wal85]. parallelism [Obe88]. Parallelizing [Pou88a]. Paranoia [Kar85]. Part [Pou89e]. Pascal [JM84, But89, JM84, JS85, Ker89, Pou88c, WH87]. Pascals [Ude89a]. Patches [JM84]. pattern [Kur89]. PC [Die89, Gla89c, ADG89a, Bar89b, Bar89e, Bir83, Che89, Han87, Lan88a, Mee89b, Mye88b, NU89, Ras89d, Smi89b, VC89b, VC89e, Wal88b, Wal88a, Wil88a, Wil88c, Won87]. PC-DOS [Che89]. PC-LAN [ADG89a]. PCs [Ano89c, MB89, Tit89, Wil88c]. Pepperoni [Sch89a]. perceptions [Daw89a]. perfect [Cra89g]. performance [Pou89c, Saa89, Sat89b]. Perils [Gru88]. Personal [BD84, Mye88a, Pou86a, Pou88d, Ste89a, Won86, Bar89h, Com89a, Kin89, MT89b, Nic89a, Obe88, Pou89c, Rob89d]. perspective [DA88, Sar89]. philosophy [Gib89]. pick [WSW89]. picture [Nic88]. pictures [Ano89e, Woo89a]. Pioneer [Byt88a]. pitfalls [Gru88]. Pixels [Kli89c]. planning [Way88a]. Plotters [DA88]. Plotting [Sch86d]. PM [Min89a]. pocket [Bar89e]. Point [Gre89d, Kar85, Wil88b, Gre88a, Gre88c]. Portable [Sha88g, Kli89b, Loe89a, Ras88c, TH89, Van89b, VC89e]. PostScript [AD88d, Egl89]. Potential [Bor87]. Power [Hom89, Pou82, Ren89, Tho89d, Ude89d, Wal88b, Wal88a, Wal88a]. PowerFlex [Hay89c]. Powerful [Kli89b, Smi83, Bar89g, TH89]. powerhouse [Nad89, SM89]. Powerlite [Sat89a]. Powers [DHM89a, DHM89b]. Practical [Wit87b]. Predicting [Sch86c]. Premium [Hol88a]. Premium/386 [Hol88a]. Prentice [Wit87b]. Prentice-Hall [Wit87b]. presentation [Ste89b, AD88c]. Press [Vos89, Ude89d]. price [Egl89]. Primer [Clo87]. Principles [Ing81]. printer [Egl89, Qui89a, Qui89b]. printers [AD88d]. priorities [Min89c]. private [Ste89a]. Probabilistic [Hin86]. problem [Mac86]. Problems [BRS80b, LM85, BRS80a, JM84, RS87b]. Processing [Byt88b, Dee85, Knu86, Obe87, Wil88c, FHMC89, Hou89, MT89a, Pou85b, Sha89b, WSW89, Wal85, WP89, Woo89b]. processor [Cra89g, Gab88, Woo88d]. Processors [Her85]. Produce [Rob87a]. Product [DA89, Ano89e]. productivity [Ras88a, Ras89c]. products [Ano89d]. Program [Dar85, JS85, Ste87, DeM88, Eug89, Gab88, Gre88d, Ros89]. programmer [Hom89]. Programmers [JS85, Ude89b]. Programming [FG84, HK85, Hoa86, Lan88b, Mye88b, Wil85, BRS80a, DHM89b, Gel88, Lin88, NU89, RS87b, Rub89b, Tho89a, Tho89h, Val89a, Weg89, Las87]. Programs [BBM80, Edg87, Pou89e, Lic89, MS88]. project [SS89, Woo88c, Woo89a]. PROLOG [Pou84a, Wei84, Clo87, Col87, Lan88c, Lan89b, Pou88a, Ste87]. Prolog-III [Col87]. promise [Ros89, Woo88c]. prompt [Mia89a]. Proofs [Way87]. Propagation [JH87]. ProPoint [Sha89c]. Protected [Gla89k]. protection [Wal88a]. protocol [Dav89a, Mee89d]. Prototyper [Val89a]. PROUST [JS85]. Provides [Phi86, Mac86]. PS [Kli89c, Smi89a, Van89c]. PS/ [Smi89a].

PS/2 [Kli89c, Van89c]. Pseudorandom [Whi84]. Public [Smi83]. Publishing [Ude89d]. pulling [Che89]. punch [Van89b]. Pushing [Eug89]. put [Tho89f]. Putting [Saf89].

Qsim [KCB89]. quantum [Ree89]. quest [Cla89]. Queue [Ras84]. quick [WN88]. QuickBASIC [Sha89a]. QuickCapture [WN88]. QuickDraw [Tho89c]. quickly [Fid88]. Quintet [Ras88d].

radio [Wll89]. Random [MSC87, WH87]. Random-Number [WH87]. Rating [Bro87]. reading [Rob89a]. reality [Daw89a]. really [Nic89b]. Reasoning [Ste87]. recognition [Kur89, Rob89a]. Record [Wei84]. redefined [Woo88b]. Redefining [Hel89b]. reference [Fid88, Mye88a]. regression [Kyd88]. Rehearsal [FG84]. Reif [Phi86, Phi87]. Rekursiv [Pou88b, Pou88e]. related [Rub89b]. relational [MB89]. release [Ren89]. rely [Dro89]. Remote [CS88, MT89a]. Remote-control [CS88]. RenderMan [Apo89]. Reproducing [BBM80]. required [Gre88c, Gre89g, Rub89b]. Resident [Edg87, Mye88b]. Resource [Ano88, Loe89b]. retrieval [All89]. retrospective [Cra89f]. Reverse [Ste85]. Review [Ano86a, Bou88, Hoe87, Ken89, Ker85, Mil84, Mye88a, Mye88b, Vos89, Wit87b, Wit87a, Ung89a]. Revisited [Gre89d, Lan88c, Min89d]. revival [Sha88b]. Revolution [Win85]. riddle [MB89]. right [BC88]. Ring [Gla89e]. RISC [HB89b, Mar84, MT89b, SM89, Way89a]. Risky [Mac89]. road [Pou84c, VC89c]. Robot [JHW86]. Robotics [Ano86a]. ROM [Ore88]. ROMs [Byt88c]. roundup [Mal88b]. Routine [WH87]. rows [Lic89]. RT [Han87]. rule [JH87, Lan89b]. Rules [TT86, Lan89b].

Safe [Woo89c, Koc89]. safety [Ano89d]. SAS [Bas84]. Save [Byt88d]. scan [Bre89]. scanners [Loe88, Tho89e, VC89b]. SCHEME [Cli88, Won87]. Science [Ken89, Vos89, Rob89d, Knu86]. Scientific [Ker85, Mun89]. screen [AD89b, Rob89c]. screens [Ras88b]. scripts [Mac86]. SE [Bar89d]. Search [Wit87a]. Secret [Dro89]. secure [Woo89c]. security [Ano89d, Dro89, Koc89, Ste89a]. Self [BBM80]. Self-Reproducing [BBM80]. Semantics [Cli88]. Semidisk [Pou82]. Separation [DHM89a, DHM89b]. sequel [VC89f]. Sequences [WH87]. server [Smi89b, vNC89]. serves [Gre89e]. Serving [AP89, vNC89]. Set [Sch86d]. seven [JM84]. Sharing [Dav89b]. sheet [Lic89]. shell [Com89b, Mia89b]. showing [Cra89e]. shows [Ros89]. SIAs [Api89]. Side [Obe88]. similar [Mac86]. simplifies [Tho89h]. simulate [Obe88]. Simulating [Hin86]. Simulation [Bro84, Ras84, Sch84, KCB89, Val89b]. Slick [Sha88d]. slides [Gro80]. SLT [Hay88a, Hol89c]. SLT/286 [Hol89c]. small [Tho89d]. Smalltalk [Cra89h, Sha88e, Ano81, Cra89c, GR81, Ing81, Tes81, Val89c]. Smalltalk-80 [Cra89c, GR81]. Smalltalk/V [Val89c]. SmethersBarnes [Val89a]. Smoothing [AU89]. Snyder [Wit87a]. Snyder-Palmer [Wit87a]. Social [Ker85]. Soft [Sha88f]. Soft-ICE [Sha88f]. Software [Dee85, Mil84, Pou82, Rob81, Sha88f, CM88, Gla89g, Lan89a, OGY^{+}89, Ras89c, Rob89b, Sha88g, Ste89b, Tho89h, Tut89, WN88, Ude89d]. Solution [Phi86, BRS80a, Mac86]. Solutions [JM84]. Solve [Ste87]. Solver [Mil84]. solving [RS87b]. Some [Gre88c, Gre89g]. sophisticated [Ung89c]. Sounds [SP89]. speak [OGY^{+}89]. Special [Ano81, Ano86b, Pou84b]. spectrum [Tho89c, Tho89e]. speech [Kur89, WH89]. speed [Che89, Gab88, Hal89a]. speeds [Tho89h]. Spell [Sha89c]. SpinRite [Gre88d]. spread [Lic89]. spreadsheet [Cra89b, Gab89a]. spreadsheets [Pea88]. Sprint [Woo88d]. SQL [Sch89b, VC89f]. Standard [Mie84, Gre81, Nic89b, Eug89]. standards [Hel89b]. Star [SIK^{+}82]. stars [ADG89a]. State [KR88, Min89e]. Step [Tho89b, Tho89h]. stops [Che89]. Storage [Nel88, AE89]. store [Van88a]. story [Obe88]. straight [Min89c]. Stratified [AS88]. Strengthening [Hal89b]. Stretching [Hay89d]. strings [Pou88c]. Structures [Han87]. style [Sha88f, Van89a]. subtle [JM84]. Suit [Woo88d]. sum [Rob87b]. Sun [Bar89g]. Sun386i [Ung88b]. Super [Las89, Nic89b]. Superbase [Bar89f]. SuperCard [Las89]. supercomputer [Cia88b, Cia88c, Cia88d]. Supercomputers [Pou86a]. SupersPort [Ung89c]. support [Bat88, NU89]. survey [HU89]. Survival [Wil88b]. SX [Hol88d, Sat89a, VC89d]. Symbolic [Way89b]. System [GR81, Hin86, Hol88c, Mye88a, Pou89b, Smi83, Wil83, But89, Dro89, Gre81, PK88, Sch89a, TS89, Tut83, Ude89c, Ung88a, VC89a, Pou89d, Yag89]. System/2 [Mye88a]. Systems [Ano86b, Hew85, Rob81, ADG89a, Bar89g, Gla89a, Kur89, Rob89a, Sha88c].

T800 [Ste88]. tablets [DA89]. Take [WSW89]. takes [Bar89d]. Taking [PKAG88]. talk [Gla89d, Sha88e]. Tandy [Van89c]. tech [Van89a, Woo88a]. Technical [Mye88a, RS87b, Ude89d]. technologies [Bre89]. Technology [Bor87, Bre89, Lan89c, PKAG88, SB88]. Ted [Byt88a]. Ten [Lic89]. terminal [CV89]. Testing [MSC87, Whi84]. tests [Pea88]. Text [All89, Han87, Knu86, Fri88, Sha88d, Ude89b]. theft [AK89]. them [JM84]. Theory [Bha87]. there [Mee88d]. thermal [Qui89a]. Thing [Wit87a]. third [HU89, LG88]. Three [Bla89, Jos87, LB81, SS80]. Three-dimensional [SS80]. thy [Gre89a]. Time [OB89]. tiny [Mun89]. titans [Coo89]. TK [Mil84]. TLC [Won85]. TLC-LISP [Won85]. today [Obe88]. together [EL88]. Token [Gla89e]. tomorrow [Obe88]. tool [BC88, Sch89c, Way88a]. Toolbox [Cha86, Opp88b]. toolkit [KCB89, Ude88, Val89b, Val89d]. Tools [Pou82, Ras88a]. TopSpeed [Nan89]. Total [Woo88e]. tour [Whi89]. Trackball [Sha89c]. transfer [Egl89, Mee89d]. Transformation [Dar85]. transistor [Ree89]. Translating [Bas84]. transparent [WP89]. Transputer [Wil88c, PK88, Pou84b, Pou88d, Wal85]. Trees [Gre89h, Gre89g]. Trees’n [Gre89i]. Trilogy [Lan88b]. TRON [SS89]. trouble [Saa89]. TRS [CC80]. TRS-80 [CC80]. true [Obe88]. Turbo [But89, Ker89, Lan88c]. Turbocharging [Pou86b]. Turing [Ken89, Mal87]. Tutor [AR85]. Tutorial [Bai85]. Twelve [Byt88c]. twist [All89, Bre89, Hol88a]. Two [Bar89g, Bar89h, Byt88c, Won86]. Typesetting [Mac86, Tut83, Hoe87].

ultimate [Die89]. Ultra [Kli89d]. UltraLite [VC89e]. Understanding [Gla89f, Loe89b]. Universe [Col87]. University [Han87]. Unix [FP89, Woo89c, Tut83, Com89a, Com89b, DO88, Gre81, Gre89c, Hoe88, Hol88b, Ker89, Lev88, Smi89b, Smi89c, TS89, Ung88b, Ung89b]. Unix/DOS [Hol88b]. Untangling [Pou88c]. upgradable [Hay89c]. upgrade [Die89]. Upscale [AD89b, Hol89d]. UPSs [ADG89b]. Use [Smi83]. User [SIK^{+}82, DHM89b, HB89a, Hoe88, Tel88]. users [Ren89]. Uses [Ste87, Wit87b]. Using [Edg87, Yan89, HK85]. utility [Eug89].

V [Val89c]. validity [Pea88]. Variations [Rob89c]. VAXes [Tit89]. Vectra [Bar89b].

Venture [Mac89]. versatile [Hay89c]. version [Che89]. versus [New88]. Very [WH87]. Very-Long-Cycle [WH87]. VGA [Kli89a, Nic89b, Wil88d]. video [Gla89h, Tho89f, Wil88d]. video-graphics [Tho89f]. viral [Gre89a]. Virtual [Kae81, Val89d]. viruses [AK89, Gre89a]. Vision [Fid88]. Visual [Ben85]. visualisation [Mun89]. VLIW [Way89a]. VMM [Hel89a]. voice [Ano89e, Sch89a]. voice/document [Sch89a]. Volumes [Byt88c]. Vose [Knu86]. Voting [Sch84]. vs [Bry89a, Gab88, Hay89a].

Wadlow [Mye88b]. Waltz [Won86]. Watcom [Hom89]. waves [Wll89]. wax [Egl89]. wax-transfer [Egl89]. way [BS88, SD88]. wealth [Dav89b]. Weighted [Sch84]. Wells [Van89a]. Wesley [Wit87a]. Where [Win88]. Which [Wat89]. white [Ras88b]. Whither [HS89, Lev88]. Will [Byt88d, Obe88]. Williams [Knu86]. Window [Pou89d, Yag89, Woo89b]. windowing [But89]. Windows [Lan89a, Val89d]. winner [Van88a]. without [Gre88a, Gre88c]. wizards [Bar89i]. Word [Byt88b, Woo88e, Cra89g, Gab88, Sha89b, Woo88d, Woo89b]. Words [Ano89e, Byt88c]. work [Saf89, Tho89f]. Working [EL88]. WorkPlus [Sha89c]. Works [Sha89c]. Workstation [Pou89b, Die89, Ung88b]. workstations [Bar89h, Com89a, Kin89, MT89b, Nic89a, Rob89d]. world [Loe88, Mia89b, MHKO89, Pas89]. worlds [Bar89h]. Worms [Ras88d]. Worth [MT89b]. wraps [PKAG88]. Wresch [Wit87b]. write [Way88a]. Writing [MS88, WA89].

X [Mee89a, Pou89d, Yag89]. X- [Mee89a]. X.25 [Sat89b]. X.400 [Mee88c]. Xenix [Gre81]. Xerox [Wad81]. XMODEM [Mee89d].

Y- [Mee89a]. years [Lic89, Rob89d].

yourself [Woo88d].

Zenith [Ude89c, Ung89c]. Zero [Way87]. Zero-Knowledge [Way87]. ZMODEM [Mee89a]. zone [Ano89d].

References

Apiki:1988:MC

[AD88a] S. Apiki and S. Diehl. 80386s for the masses (80386-based clones). Byte Magazine, 13(10):164–176, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1988:BAG

[AD88b] S. Apiki and S. Diehl. Benchmarks at a glance (microcomputers). Byte Magazine, 13(13):236–238, 240, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1988:PML

[AD88c] S. Apiki and S. Diehl. Presentation Manager and LAN Manager (OS/2). Byte Magazine, 13(10):157– 159, October 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1988:PPC

[AD88d] Steve Apiki and Stan Diehl. Post-Script printers come of age (buyer’s guide). Byte Magazine, 13(9):164– 178, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:BNE

[AD89a] S. Apiki and S. Diehl. 4800 bits, no errors (modems). Byte Magazine, 14(6):162–172, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:UML

[AD89b] S. Apiki and S. Diehl. Upscale monitors (large-screen monitors). Byte Magazine, 14(3):162–172, 174, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:BNS

[ADG89a] S. Apiki, S. Diehl, and R. Grehan. Battle of the network stars (PC-LAN operating systems). Byte Magazine, 14(7):154, 156, 158, 160, 162, 164, 166, 168–170, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:CBB

[ADG89b] S. Apiki, S. Diehl, and R. Grehan. Curing the brownout blues (UPSs). Byte Magazine, 14(4):162– 176, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:OOS

[AE89] S. Apiki and H. Eglowstein. The optical option (storage). Byte Magazine, 14(10):160–174, October 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:BBG

[AEG89] S. Apiki, H. Eglowstein, and R. Grehan. The brains behind the graphics. Byte Magazine, 14(12):178–198, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Adney:1989:DBV

[AK89] W. M. Adney and D. E. Kavanagh. The data bandits (viruses, theft, and accidental destruction). Byte Magazine, 14(1):267–270, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1989:TRT

[All89] D. Allen. Text retrieval with a twist. Byte Magazine, 14(7):201–202, 204, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Amsterdam:1986:DCH

[Ams86] Jonathan Amsterdam. Data compression with Huffman coding. Byte Magazine, 11(5):99–108, May 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1981:SIS

[Ano81] Anonymous. Special issue on Smalltalk. Byte Magazine, 6(8):??, August 1981. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1986:RIR

[Ano86a] Anonymous. Review of Introduction to Robotics by Arthur J. Critchlow. Byte Magazine, 11(1):57–60, January 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1986:SIO

[Ano86b] Anonymous. Special issue on object-oriented systems. Byte Magazine, 11(8):??, August 1986. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1988:LRG

[Ano88] Anonymous. LISP resource guide.

Byte Magazine, 13(2):236, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1989:DFD

[Ano89a] Anonymous. The data file (DBMS).

Byte Magazine, 14(9):291–293, September 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1989:MCC

[Ano89b] Anonymous. Making a case for CASE (buyer’s guide). Byte Magazine, 14(13):154–171, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1989:MMF

[Ano89c] Anonymous. Megahertz madness (fast PCs). Byte Magazine, 14(11): 13–18, 20, 22, 26, 28, 30–32, 34, 36, 38, 42, 44, 46, 48, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1989:SZS

[Ano89d] Anonymous. The safety zone (security products for microcomputers). Byte Magazine, 14(6):290–291, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1989:WPV

[Ano89e] Anonymous. Words and pictures (voice and image product guide). Byte Magazine, 14(13):320–321, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anderson:1989:SMM

[AP89] B. T. Anderson and M. A. Puhnaty. Serving many masters (bus masters). Byte Magazine, 14(11): 131–132, 134, 136, 138, 140, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:MCS

[Api89] S. Apiki. 32.5 MHz and climbing SIAs 386/32. Byte Magazine, 14(4): 106–109, April 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apodaca:1989:RI

[Apo89] T. Apodaca. The RenderMan interface. Byte Magazine, 14(4):167– 176, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anderson:1985:LTI

[AR85] J. R. Anderson and B. J. Reiser. The LISP tutor: It approaches the effectiveness of a human tutor. Byte Magazine, 10(4):159–175, April 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Abelson:1988:LLS

[AS88] Harold Abelson and Gerald Jay Sussman. LISP: A language for stratified design. Byte Magazine, 13 (2):207–218, February 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apiki:1989:SCO

[AU89] S. Apiki and J. Udell. Smoothing out C (optimizing C compilers). Byte Magazine, 14(2):170–186, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bailey:1985:HT

[Bai85] R. Bailey. A Hope tutorial. Byte Magazine, 10(8):235–255, August 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:ANC

[Bar89a] N. Baran. Apple’s new compact Mac IIx. Byte Magazine, 14(5):117–118, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:EAH

[Bar89b] N. Baran. EISA arrives (HP Vectra 486 PC). Byte Magazine, 14(12):93– 98, November 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:GC

[Bar89c] N. Baran. A great communicator.

Byte Magazine, 14(1):195–196,198, January 1989. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:MST

[Bar89d] N. Baran. The Mac SE takes off. Byte Magazine, 14(2):113–116, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:PYP

[Bar89e] N. Baran. A PC in your pocket. Byte Magazine, 14(12):115–117, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:SGO

[Bar89f] N. Baran. Superbase 4 (graphics oriented DBMs). Byte Magazine, 14 (3):221–222, 224, March 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:TPS

[Bar89g] N. Baran. Two powerful systems from Sun. Byte Magazine, 14(5): 108–112, May 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:TWC

[Bar89h] N. Baran. Two worlds converge (personal workstations). Byte Magazine, 14(2):229–233, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barron:1989:WML

[Bar89i] J. J. Barron. The wizards of the Media Lab. Byte Magazine, 14(13):

353–360, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bass:1984:TSL

[Bas84] J. Bass. Translating the SAS language into BASIC. Byte Magazine, 9 (1):417–434, January 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Batory:1988:ADA

[Bat88] C. J. Batory. Adding dimensions (array support in C). Byte Magazine, 13(12):419–424, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Burger:1980:SRP

[BBM80] J. Burger, D. Brill, and F. Machi.

Self-reproducing programs. Byte Magazine, 5(??):??, ?? 1980. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Begeman:1988:RTJ

[BC88] M. L. Begeman and J. Conklin.

The right tool for the job (hypertext). Byte Magazine, 13(10):255– 266, October 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bortz:1984:LIP

[BD84] J. Bortz and J. Diaman. LISP for the IBM personal computer. Byte Magazine, 9(??):281–291, July 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Benzon:1985:VMM

[Ben85] B. Benzon. The visual mind and the Macintosh. Byte Magazine, 10(??): 113–130, January 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Betz:1988:EL

[Bet88] D. Betz. Embedded languages. Byte Magazine, 13(12):409–416, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bharath:1987:IT

[Bha87] ?. Bharath. Information theory. Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Birenbaum:1983:IPM

[Bir83] Larry Birenbaum. The IBM PC meets Ethernet. Byte Magazine, 7 (??):272–280, November 1983. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Brown:1980:FCG

[BL80] B. E. Brown and S. Levine. The future of computer graphics. Byte Magazine, 5(11):22–28, November 1980. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Blaszczak:1989:TAM

[Bla89] M. Blaszczak. Three assemblers for MS-DOS. Byte Magazine, 14(2): 205–209, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baran:1989:EL

[BN89] N. Baran and M. E. Nadeau. Everexpanding laptops. Byte Magazine, 14(8):90–96, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bork:1987:PIT

[Bor87] ?. Bork. The potential for interactive technology. Byte Magazine, 12 (??):??, ?? 1987. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bourn:1988:RDM

[Bou88] ?. Bourn. Review of Duncan, The MS-DOS Encyclopedia (1988). Byte Magazine, 13(??):??, ?? 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bretzmann:1989:NTO

[Bre89] J. Bretzmann. A new twist on an old technology helical-scan technologies. Byte Magazine, 14(12):380– 388, November 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bronson:1984:CSW

[Bro84] R. Bronson. Computer simulation:

What it is and how it’s done. Byte Magazine, 9(3):95–102, March 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Brookshire:1987:RIC

[Bro87] R. G. Brookshire. Rating the IBM compatibles. Byte Magazine, 12(12): 193–6, 198, December 1987. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Berresford:1980:KAN

[BRS80a] G. C. Berresford, A. M. Rockett, and J. C. Stevenson. Khachiyan’s algorithm. I. A new solution to linear programming problems. Byte Magazine, 5(8):198–208, August 1980. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Berresford:1980:KAP

[BRS80b] G. C. Berresford, A. M. Rockett, and J. C. Stevenson. Khachiyan’s algorithm. II. problems with the algorithm. Byte Magazine, 5(9): 242–255, September 1980. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryce:1989:FVM

[Bry89a] J. Y. Bryce. Fiber vs. metal-mix, match, and choose between fiberoptic and metallic cables. Byte Magazine, 14(1):253–258, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bryce:1989:GPN

[Bry89b] J. Y. Bryce. Growing pains (network OS). Byte Magazine, 14(8):135–136, 138, August 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barnsley:1988:BWC

[BS88] M. F. Barnsley and A. D. Sloan. A better way to compress images. Byte Magazine, 13(??):215–223, January 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Butler:1989:TPW

[But89] C. J. Butler. Turbo Pascal windowing system. Byte Magazine, 14(2): 283–291, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Betz:1988:IL

[BW88] David Betz and Gregg Williams. Introduction LISP. Byte Magazine, 13(2):206, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Byte:1988:HHP

[Byt88a] Byte staff. Hyperquotes from hypertext pioneer Ted Nelson. Byte Magazine, 13(2):14, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Nanobytes Department.

Byte:1988:HWP

[Byt88b] Byte staff. Hypertext word processing. Byte Magazine, 13(2):84, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). What’s New Section (product announcements).

Byte:1988:TVE

[Byt88c] Byte staff. Twelve volumes of English words now on two CD-ROMs. Byte Magazine, 13(2):18, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Microbytes department.

Byte:1988:WHS

[Byt88d] Byte staff. Will hypertext save mankind? Byte Magazine, 13(2):

14, February 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Microbytes department.

Cohen:1980:GIG

[CC80] D. K. Cohen and D. Crowe. A general interpolating graphics package for the TRS-80. Byte Magazine, 5 (11):296–310, November 1980. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chang:1986:AT

[Cha86] Y. F. Chang. The ATOMCC toolbox. Byte Magazine, 11(4):215– 224, April 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chen:1989:PPS

[Che89] F. Chen. PC-DOS: pulling out the stops (version 4.0 features to speed up file access). Byte Magazine, 14 (6):294–299, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1984:L

[Cia84] S. A. Ciarcia. The lis’ner 1000. Byte Magazine, 9(??):111–124, November 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1986:BHD

[Cia86] Steve Ciarcia. Build a hardware data encryptor. Byte Magazine, 11 (??):??, ?? 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:CCCb

[Cia88a] S. Ciarcia. Ciarcia’s circuit cellar. 2. why microcontrollers? Byte Magazine, 13(9):303–312, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:CCCc

[Cia88b] S. Ciarcia. Ciarcia’s circuit cellar. 3 A supercomputer. Byte Magazine, 13(13):327–339, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:Sa

[Cia88c] S. Ciarcia. A supercomputer. 1.

Byte Magazine, 13(10):283–291, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:Sb

[Cia88d] S. Ciarcia. A supercomputer.

2. Byte Magazine, 13(12):399–406, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clarkson:1989:QMC

[Cla89] M. A. Clarkson. The quest for the molecular computer. Byte Magazine, 14(5):268–273, May 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clinger:1988:SS

[Cli88] William Clinger. Semantics of SCHEME. Byte Magazine, 13(2): 221–227, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clocksin:1987:PP

[Clo87] W. Clocksin. A Prolog primer. Byte Magazine, 12(??):147–158, August 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crowston:1988:ISA

[CM88] K. Crowston and T. W. Malone. Intelligent software agents (groupware). Byte Magazine, 13(13):267– 272, December 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Colmerauer:1987:OPI

[Col87] A. Colmerauer. Opening the Prolog-III universe. Byte Magazine, 12 (9):177–182, August 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Comeau:1989:NUP

[Com89a] G. Comeau. Networking with Unix (personal workstations). Byte Magazine, 14(2):265–267, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Comeau:1989:US

[Com89b] G. Comeau. The Unix shell. Byte Magazine, 14(9):315–316, 318–321, September 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1984:EE

[Coo84] P. R. Cook. Electronic encyclopedias. Byte Magazine, 9(??):151–167, July 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1989:CGT

[Coo89] R. Cook. Clash of the graphics titans. Byte Magazine, 14(11):143– 144, 146–148, 150, 152, 154, 156, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1988:MBC

[Cra88] D. Crabb. To Mac and back (computer links). Byte Magazine, 13(11): 67–70, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:DEM

[Cra89a] D. Crabb. DAT drive eases Mac backups. Byte Magazine, 14(12): 225–230, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:FS

[Cra89b] D. Crabb. The flying spreadsheet.

Byte Magazine, 14(7):207–208, 210, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:HSM

[Cra89c] D. Crabb. Hooked on Smalltalk-80 for the MAC. Byte Magazine, 14(1):143, 145–146, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:MGF

[Cra89d] D. Crabb. Mac goes fax. Byte Magazine, 14(5):208C–D, F, H, May 1989.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:MIS

[Cra89e] D. Crabb. The Mac interface: showing its age. Byte Magazine, 14(6): 235–237, June 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:MR

[Cra89f] D. Crabb. A Macintosh retrospective. Byte Magazine, 14(3):143–144, 146, March 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:PWP

[Cra89g] D. Crabb. A perfect word processor at last? Byte Magazine, 14(5): 157–158, May 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Crabb:1989:SCB

[Cra89h] D. Crabb. Smalltalk can be cheap. Byte Magazine, 14(4):141– 146, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cook:1988:RC

[CS88] R. Cook and P. Schauble. Remotecontrol communications. Byte Magazine, 13(12):249–252, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Catchings:1989:LTA

[CV89] B. Catchings and M. L. Van Name. The LAN terminal alternative. Byte Magazine, 14(12):211–214, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1988:PPB

[DA88] S. Diehl and S. Apiki. Plotters in perspective (buyer’s guide). Byte Magazine, 13(13):162–180, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1989:PFD

[DA89] S. Diehl and S. Apiki. Product focus: digitizing tablets: graphic details. Byte Magazine, 14(1):162–166, 168–172, 174, January 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Darlington:1985:PT

[Dar85] J. Darlington. Program transformation. Byte Magazine, 10(8):201–218, August 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1989:LCA

[Dav89a] R. Davis. A logical choice (APPC or LU 6.2 protocol). Byte Magazine, 14 (1):309–315, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Davis:1989:SWD

[Dav89b] R. Davis. Sharing the wealth (distributed databases). Byte Magazine, 14(9):267–270, 272–274, September 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1989:CPR

[Daw89a] B. M. Dawson. Changing perceptions of reality (image editing). Byte Magazine, 14(13):293–304, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dawson:1989:FMO

[Daw89b] J. Dawson. A family of models (object-oriented databases). Byte Magazine, 14(9):277–278, 280–282, 284, 286, September 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Deering:1985:AAH

[Dee85] M. Deering. Architectures for AI: Hardware and software for efficient processing. Byte Magazine, 10(4): 193–206, April 1985. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

DeMaria:1988:MID

[DeM88] R. DeMaria. Macdraw II (drawing program). Byte Magazine, 13(13): 231–234, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dodani:1989:SP

[DHM89a] ?. Dodani, ?. Hughes, and ?. Moshell. Separation of powers. Byte Magazine, 14(??):??, ?? 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dodani:1989:SPU

[DHM89b] M. H. Dodani, C. E. Hughes, and J. M. Moshell. Separation of powers (user interfaces and object-oriented programming). Byte Magazine, 14 (3):255–256, 258, 260–262, 270–271, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Diehl:1989:UUI

[Die89] S. Diehl. The ultimate upgrade (IBM PC AT upgrade to 80386 based workstation). Byte Magazine, 14(6):313–319, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dougherty:1988:MU

[DO88] D. Dougherty and T. O’Reilly. DOS meets Unix. Byte Magazine, 13(11): 117–126, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dror:1989:SCG

[Dro89] Asael Dror. Secret codes (any good data security system must rely on encryption). Byte Magazine, 14(6): 267–270, June 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Duff:1987:DEL

[Duf87] Charles B. Duff. Designing and efficient language. Byte Magazine, 12 (??):211–224, August 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Edginton:1987:IMR

[Edg87] ?. Edginton. Installing memoryresident programs with C: Using C programs to extend DOS. Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eglowstein:1989:WPC

[Egl89] H. Eglowstein. What price color PostScript? (wax-transfer printer). Byte Magazine, 14(13):229–232, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Engelbart:1988:WTG

[EL88] D. Engelbart and H. Lehtman. Working together (groupwise). Byte Magazine, 13(13):245–252, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eisenbach:1985:DLO

[ES85] S. Eisenbach and C. Sadler. Declarative languages: An overview. Byte Magazine, 10(8):181–200, August 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Eugenides:1989:PSF

[Eug89] J. Eugenides. Pushing Standard File to the limit (utility program). Byte Magazine, 14(6):225–231, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Feldman:1985:C

[Fel85] J. A. Feldman. Connections. Byte Magazine, 10(??):277–284, April 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Finzer:1984:PR

[FG84] W. Finzer and L. Gould. Programming by rehearsal. Byte Magazine, 9(6):187–210, June 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fox:1989:HPP

[FHMC89] G. C. Fox, A. W. Ho, P. Messina, and T. Cole. Hands-on parallel processing. Byte Magazine, 14(10):287– 293, October 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fichtelman:1985:EM

[Fic85] M. Fichtelman. The expert mechanic. Byte Magazine, 10(6):205– 216, June 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiderio:1988:GVH

[Fid88] J. Fiderio. A grand vision–hypertext mimics the brain’s ability to access information quickly and intuitively by reference. Byte Magazine, 13(10): 237–244, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fiedler:1989:FIU

[FP89] D. Fiedler and J. Paul. Future imperfect (unix). Byte Magazine, 14(5):227–228, 230–234, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Frey:1986:BMC

[Fre86] P. W. Frey. A bit-mapped classifier.

Byte Magazine, 11(??):??, ?? 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Frisse:1988:TH

[Fri88] M. Frisse. From text to hypertext. Byte Magazine, 13(10):247– 253, October 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fried:1989:ONC

[Fri89] S. Fried. Optimizing numeric coprocessing. Byte Magazine, 14(11):221– 224, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gabaldon:1988:FVS

[Gab88] D. Gabaldon. Features vs. speed (FullWrite word processor program). Byte Magazine, 13(12):241– 246, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gabaldon:1989:FIS

[Gab89a] D. Gabaldon. Full Impact (spreadsheet package). Byte Magazine, 14 (2):211–214, February 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gabaldon:1989:MPD

[Gab89b] D. Gabaldon. Move over, Page-Maker (DTP package). Byte Magazine, 14(13):217–220, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gelernter:1988:GJD

[Gel88] David Gelernter. Getting the job done (linda, parallel programming language). Byte Magazine, 13(12): 301–308, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gibson:1989:CP

[Gib89] M. L. Gibson. The CASE philosophy. Byte Magazine, 14(4):209– 218, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Giguere:1989:EOD

[Gig89] E. Giguere. Electronic Oxford (dictionary computerisation). Byte Magazine, 14(13):371–374, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gilliland:1988:OC

[Gil88] J. Gilliland. OS/2 communications. Byte Magazine, 13(11):153– 166, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:CMS

[Gla89a] B. Glass. Caching in on memory systems. Byte Magazine, 14(3):281– 285, March 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:HDI

[Gla89b] B. Glass. Hard disk interfaces. Byte Magazine, 14(2):293–297, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:IPB

[Gla89c] B. Glass. The IBM PC BIOS. Byte Magazine, 14(4):303–310, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:MAT

[Gla89d] B. Glass. Making applications talk (CAS). Byte Magazine, 14(1):155– 158, January 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:TR

[Gla89e] B. Glass. The token ring. Byte Magazine, 14(1):363–368, 370, 372, 374, 376, January 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:UN

[Gla89f] B. Glass. Understanding NetBIOS. Byte Magazine, 14(1):301–306, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:HDM

[Gla89g] K. B. Glass. Hard disc maintenance software. Byte Magazine, 14 (8):265–268, 270, 272, 274, 276, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:DVI

[Gla89h] L. B. Glass. Digital video interactive. Byte Magazine, 14(5):283–289, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:LAE

[Gla89i] L. B. Glass. The light at the end of the LAN. Byte Magazine, 14(7):

269–274, July 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:MMM

[Gla89j] L. B. Glass. Modern modem methods. Byte Magazine, 14(6):321–326, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Glass:1989:PM

[Gla89k] L. B. Glass. Protected mode. Byte Magazine, 14(13):377–384, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Goldberg:1981:SS

[GR81] A. Goldberg and D. Robson. The Smalltalk-80 system. Byte Magazine, 6(8):36–48, August 1981. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Graef:1989:GF

[Gra89] G. L. Graef. Graphics formats. Byte Magazine, 14(9):305–306, 308–310, September 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Greenberg:1981:UOS

[Gre81] Robert B. Greenberg. The Unix operating system and the Xenix standard operating environment. Byte Magazine, 6(6):248–264, June 1981. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1988:FPC

[Gre88a] R. Grehan. Floating-point without a coprocessor. Byte Magazine, 13(9):

313–319, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1988:OO

[Gre88b] R. Grehan. An overview of overlays. Byte Magazine, 13(13):341–348, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1988:SAR

[Gre88c] R. Grehan. Some assembly required. 2. Floating-point without a coprocessor. Byte Magazine, 13(10):293– 298, October 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1988:SHD

[Gre88d] R. Grehan. Spinrite (hard disc diagnostic and maintenance program). Byte Magazine, 13(12):237– 239, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Greenberg:1989:KTV

[Gre89a] Ross M. Greenberg. Know thy viral enemy (computer viruses). Byte Magazine, 14(6):275–280, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:DA

[Gre89b] R. Grehan. Directory assistance. 1. Byte Magazine, 14(5):291–294, 296– 298, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:DAM

[Gre89c] R. Grehan. Directory assistance. 2.

(MS-DOS, Unix, Macintosh HFs). Byte Magazine, 14(6):327–336, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:FPR

[Gre89d] R. Grehan. Floating-point revisited. Byte Magazine, 14(4):311– 318, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:IMS

[Gre89e] R. Grehan. If memory serves … (memory management). Byte Magazine, 14(8):279–280, 282, 284, 286, 337, August 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:IKI

[Gre89f] R. Grehan. Input: keep it clean. Byte Magazine, 14(13):387– 392, 442, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:SAR

[Gre89g] R. Grehan. Some assembly required.

2. Trees ’n keys. Byte Magazine, 14(2):301–307, February 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:TKI

[Gre89h] R. Grehan. Trees ’n keys. III. Byte Magazine, 14(3):287–288, 290, 292– 293, March 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1989:TK

[Gre89i] R. Grehan. Trees’n keys. 1. Byte Magazine, 14(1):379–380, 384, 386– 390, January 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grogono:1980:GCS

[Gro80] A. W. Grogono. Graphic color slides. Byte Magazine, 5(11):126– 144, November 1980. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grudin:1988:PPG

[Gru88] J. Grudin. Perils and pitfalls (groupware). Byte Magazine, 13(13):261– 264, December 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1988:BBL

[GT88] R. Grehan and T. Thompson. Borland beefs up its languages. Byte Magazine, 13(10):151–154, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halliday:1989:INS

[Hal89a] C. Halliday. IBM’s new speed king. Byte Magazine, 14(7):173–176, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halliday:1989:SL

[Hal89b] C. Halliday. Strengthening the lineup. Byte Magazine, 14(1):179– 184, 186, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hansen:1987:DSB

[Han87] ?. Hansen. Data structures in a bitmapped text editor: How Carnegie-Mellon University displays text on the IBM RT PC. Byte Magazine, 12(??):??, ?? 1987. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1988:ALL

[Hay88a] F. Hayes. At long last, laptop (compaq 286 SLT laptop). Byte Magazine, 13(13):107–110, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1988:CC

[Hay88b] F. Hayes. The crossbar connection. Byte Magazine, 13(??):278– 279, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1989:BCV

[Hay89a] F. Hayes. Battle of the chips (80286 vs 80386). Byte Magazine, 14(3): 274–279, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1989:IC

[Hay89b] F. Hayes. Intel’s Cray-on-a-chip.

Byte Magazine, 14(5):113–114, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1989:PVU

[Hay89c] F. Hayes. Powerflex: a versatile, upgradable AT clone. Byte Magazine, 14(12):110–113, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1989:SL

[Hay89d] F. Hayes. Stretching DOS to the limit. Byte Magazine, 14(11):79–80, 82, 84, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1989:GGG

[HB89a] F. Hayes and N. Baran. A guide to GUIs (graphical user interfaces). Byte Magazine, 14(7):250–257, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hunter:1989:AR

[HB89b] C. Hunter and J. Banning. DOS ar RISC. Byte Magazine, 14(12):412– 418, November 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1989:BMB

[Hel89a] M. Heller. Breaking the memory barrier with 386 mod VMM. Byte Magazine, 14(7):187–188, 190, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heller:1989:RS

[Hel89b] M. Heller. Redefining the standards. Byte Magazine, 14(11):56–58, 60– 64, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hershey:1985:IP

[Her85] W. Hershey. Idea processors. Byte Magazine, 10(??):??, June 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hershey:1987:G

[Her87] William Hershey. Guide. Byte Magazine, 12(10):244–246, October 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Software Review.

Hewitt:1985:COS

[Hew85] Carl Hewitt. The challenge of open systems. Byte Magazine, 10(4):223– 242, April 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hinton:1985:LPN

[Hin85] G. E. Hinton. Learning in parallel networks. Byte Magazine, 10(??): 265–273, April 1985. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hinton:1986:LPN

[Hin86] G. E. Hinton. Learning in parallel networks: Simulating learning in a probabilistic system. Byte Magazine, 10(4):265–273, April 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Harrison:1985:FPU

[HK85] P. G. Harrison and H. Khoshnevisan. Functional programming using FP. Byte Magazine, 10(8):

219–234, August 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hayes:1989:LF

[HNM89] F. Hayes, M. E. Nadeau, and S. Miastkowski. Laptops forever. Byte Magazine, 14(13):93–96, 98, 100, 102–103, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hoare:1986:MP

[Hoa86] C. A. R. Hoare. Mathematics of programming. Byte Magazine, 10 (8):115–149, August 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hoenig:1987:RKC

[Hoe87] Alan Hoenig. Review of Knuth’s Computers & Typesetting. Byte Magazine, 12(??):??, ?? 1987. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hoeber:1988:FFO

[Hoe88] T. Hoeber. Face to face with Open Look (Unix user interface). Byte Magazine, 13(13):286–296, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1988:TAP

[Hol88a] J. Holtzman. An 80386 with a twist (AST Premium/386). Byte Magazine, 13(9):195–200, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1988:MCU

[Hol88b] J. Holtzman. Merge 386 (combined Unix/DOS environment). Byte Magazine, 13(13):207–212, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1988:NAD

[Hol88c] J. Holtzman. A nimble AT (Dell System 220). Byte Magazine, 13 (13):193–196, December 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1988:SAC

[Hol88d] J. Holtzman. SX appeal (compaq 386s microcomputers). Byte Magazine, 13(12):197–202, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1989:AFD

[Hol89a] J. Holtzman. Advanced floppy disk drive controllers. Byte Magazine, 14 (3):191–195, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1989:ECM

[Hol89b] J. Holtzman. EMS with a cache (memory board). Byte Magazine, 14 (7):181–182, 184, July 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1989:LDC

[Hol89c] J. Holtzman. Laptop dilemma Compaq SLT/286 and Ogivar 286. Byte Magazine, 14(3):177–180, 182, 184, 186, 188–189, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtzman:1989:UAA

[Hol89d] J. Holtzman. Upscale Acer (Acer 1100/33, micro). Byte Magazine, 14 (13):185–186, December 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hommel:1989:PPW

[Hom89] F. Hommel. Power to the programmer (watcom C 386). Byte Magazine, 14(13):199–202, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hough:1989:POD

[Hou89] D. Hough. The paperless office (distributed processing). Byte Magazine, 14(7):241–242, 244–246, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Humphrey:1989:WM

[HS89] J. H. Humphrey and G. S. Smock. Whither the modem? Byte Magazine, 14(1):281–283, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Holtz:1989:TDM

[HU89] B. Holtz and J. Udell. The third dimension (3D modelling packages, survey). Byte Magazine, 14(5):178– 180, 182–186, 188–190, 192, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ingalls:1981:DPB

[Ing81] D. H. H. Ingalls. Design principles behind Smalltalk. Byte Magazine, 6 (8):286–298, August 1981. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Jones:1987:BPG

[JH87] William P. Jones and Josiah Hoskins. Back-propagation — a generalized delta learning rule. Byte Magazine, 12(10):155–162, October 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hamel:1986:ARN

[JHW86] W. Jorgensen, C. Hamel, and C. Weisbin. Autonomous robot navigation. Byte Magazine, 10(1):223– 235, January 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Johnson:1984:PDF

[JM84] Marc C. Johnson and Allen Munro.

Pascal’s design flaws: Modula-2 solutions and Pascal patches: A description of seven subtle problems with pascal, and a look at how modula-2 avoids them. Byte Magazine, 9(3):371–372, 374, 376, 378, 380, 382, 384, 387– 388, March 1984. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). URL https:

//ia800308.us.archive.org/20/
items/byte-magazine-1984-03/
1984_03_BYTE_09-03_Simulation.
pdf.

Jones:1989:WCC

[Jon89] T. C. Jones. Why choose CASE? Byte Magazine, 14(13):80IS3–4, 6, 8, 10, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Josin:1987:NNH

[Jos87] G. Josin. Neural-network heuristics: Three heuristic algorithms that learn from experience. Byte Magazine, 12(10):177–192, October 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Johnson:1985:PLP

[JS85] W. Lewis Johnson and Elliot Soloway. PROUST. the LISP program automatically debugs the efforts novice Pascal programmers. Byte Magazine, 10(4):179– 190, April 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kaehler:1981:VMO

[Kae81] Ted Kaehler. Virtual memory for an object-oriented language. Byte Magazine, 6(8):378–387, August 1981. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). The special issue on Smalltalk has a very informative tutorial.

Karpinski:1985:PFP

[Kar85] R. Karpinski. Paranoia: A floatingpoint benchmark. Byte Magazine, 10(2):223–235, February 1985. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kimbrell:1989:QST

[KCB89] R. E. Kimbrell, L. Correll, and R. Bass. The Qsim simulation toolkit. Byte Magazine, 14(7):259–266, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kenner:1989:RDT

[Ken89] ?. Kenner. Review of Dewdney, The Turing Omnibus: 61 Excursions in Computer Science (1989). Byte Magazine, 14(??):??, ?? 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kerber:1985:RCC

[Ker85] Kenneth W. Kerber. Review of computer culture: The scientific, intellectual and social impact of the computer. Byte Magazine, 10(10):57–58, October 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kerr:1989:UFT

[Ker89] J. Kerr. Unix filenames for Turbo Pascal. Byte Magazine, 14(11):185– 186, 188, 190, 192, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kindel:1989:HFF

[Kin89] B. Kindel. How fast is fast? (personal workstations). Byte Magazine, 14(2):251–254, February 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1989:DV

[Kli89a] B. D. Kliewer. Debunking 16-bit VGA. Byte Magazine, 14(6):195– 199, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1989:HPP

[Kli89b] B. D. Kliewer. HOOPS: Powerful portable 3-D graphics. Byte Magazine, 14(7):193–194, 196, 198, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1989:PMA

[Kli89c] B. D. Kliewer. Pixels on the march (8514/A and Artist 10 MC graphics coprocessor boards for the IBM PS/2). Byte Magazine, 14(1):201– 202, 204–208, January 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kliewer:1989:UGG

[Kli89d] B. D. Kliewer. Ultra Graphics (graphics coprocessor board). Byte Magazine, 14(8):167–169, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Knuth:1986:TPC

[Knu86] Donald E. Knuth. Text processing: Computer Science considerations (interview by G. Michael Vose and Gregg Williams). Byte Magazine, 11(2):169–172, February 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kochanski:1989:HSI

[Koc89] Martin Kochanski. How safe is it? (computer security). Byte Magazine, 14(6):257–264, June 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kosko:1987:CAM

[Kos87] B. Kosko. Constructing an associative memory. Byte Magazine, 12 (??):137–144, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kernighan:1988:SC

[KR88] Brian W. Kernighan and Dennis M.

Ritchie. The state of C. Byte Magazine, 13(??):??, ?? 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kurzweil:1989:BPR

[Kur89] R. Kurzweil. Beyond pattern recognition (speech recognition systems). Byte Magazine, 14(13):277–288, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kyd:1988:MRE

[Kyd88] C. W. Kyd. Multiple regression with Excel. Byte Magazine, 13(12):363– 373, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1988:LP

[Lan88a] A. Lane. Lint for the PC. Byte Magazine, 13(12):229–234, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1988:TNA

[Lan88b] A. Lane. Trilogy: A new approach to logic programming. Byte Magazine, 13(??):145–151, March 1988.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1988:TPR

[Lan88c] A. Lane. Turbo Prolog revisited. Byte Magazine, 13(10):209–212, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1989:DMW

[Lan89a] A. Lane. Domesticating Microsoft Windows (software package). Byte Magazine, 14(6):205–207, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1989:EDR

[Lan89b] A. Lane. An end to duelling rules (rule base checking in Prolog). Byte Magazine, 14(6):303–308, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lane:1989:MAH

[Lan89c] A. Lane. MicroExplorer in action! (hybrid Macintosh/Lisp technology). Byte Magazine, 14(12): 247–249, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lassez:1987:CLP

[Las87] Catherine Lassez. Constraint Logic Programming. Byte Magazine, 12 (??):171–176, August 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lasky:1989:HSS

[Las89] R. D. Lasky. How super is Super-Card? Byte Magazine, 14(10):217– 220, October 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Levitan:1981:TML

[LB81] S. P. Levitan and J. G. Bonar. Three microcomputer LISP’s. Byte Magazine, 6(??):??, September 1981. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Levitt:1988:WIU

[Lev88] J. Levitt. Whither IBM and Unix?

Byte Magazine, 13(11):109–110,114, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Little:1988:TDC

[LG88] M. J. Little and J. Grinberg.

The third dimension (3-D computer). Byte Magazine, 13(12):311– 319, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Libes:1985:B

[Lib85] Sol Libes. Bytelines. Byte Magazine, 10(9):420, September 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Licklider:1989:TYR

[Lic89] T. R. Licklider. Ten years of rows and columns (spread sheet programs). Byte Magazine, 14(13):324– 331, December 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Linowes:1988:IAO

[Lin88] ?. Linowes. It’s an attitude (objectoriented programming in C). Byte Magazine, 13(??):??, ?? 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Larimore:1985:POD

[LM85] W. E. Larimore and R. K. Mehra. Problems of overfitting data. Byte Magazine, 10(??):167–180, October 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loeb:1988:BOW

[Loe88] L. H. Loeb. Bringing the outside world into a Macintosh (scanners). Byte Magazine, 13(10):194– 199, October 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loeb:1989:PCM

[Loe89a] L. H. Loeb. A portable companion for the Macintosh (laptop computer). Byte Magazine, 14(6):241– 244, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Loeb:1989:UMR

[Loe89b] L. H. Loeb. Understanding the Macintosh resource editor. Byte Magazine, 14(7):88IS/49–52, 88IS/54, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

MacKay:1986:TPS

[Mac86] Pierre A. MacKay. Typesetting problem scripts. computer typesetting provides a solution for Arabic and similar scripts. Byte Magazine, 11(2):201–218, February 1986. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Macaluso:1989:RBI

[Mac89] P. Macaluso. A risky business — an introduction to Monte Carlo venture analysis. Byte Magazine, 14 (or 9?) (3):179–192, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Malitz:1987:TM

[Mal87] I. Malitz. The Turing machine. Byte Magazine, 12(11):345–357, November 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Malloy:1988:

[Mal88a] R. Malloy. DOS 4.0. Byte Magazine, 13(11):75–78, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Malloy:1988:MBR

[Mal88b] R. Malloy. Memory board roundup (buyer’s guide). Byte Magazine, 13 (11):81–85, 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Malloy:1989:FE

[Mal89] R. Malloy. The fastest 80386s ever?

Byte Magazine, 14(6):109–112, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Markoff:1984:RC

[Mar84] J. Markoff. RISC chips. Byte Magazine, 9(11):191–206, November 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Margulis:1989:I

[Mar89] N. Margulis. The Intel 80860. Byte Magazine, 14(13):333–340, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1989:PNE

[MB89] S. Miastkowski and N. Baran. Paradox 3: neither enigma nor riddle (relational DBMS for IBM PCs). Byte Magazine, 14(2):109–111, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McClure:1989:CE

[McC89] C. McClure. The CASE experience. Byte Magazine, 14(4):235– 236, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1988:BBB

[Mee88a] B. N. Meeks. Beefed-up bulletin boards. Byte Magazine, 13(11):45– 48, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1988:FBF

[Mee88b] B. N. Meeks. Fax board faire (buyer’s guide). Byte Magazine, 13 (9):203–210, September 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1988:XG

[Mee88c] B. N. Meeks. X.400 grows up. Byte Magazine, 13(13):149–152, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1988:YCG

[Mee88d] B. N. Meeks. You can’t get there from here-or can you? (gateways). Byte Magazine, 13(12):153– 155, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1989:AXZ

[Mee89a] B. N. Meeks. The ABCs of X-, Y- , and ZMODEM. Byte Magazine, 14(2):163–166, February 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1989:DPC

[Mee89b] B. N. Meeks. Dialing up 1990 (PC communications). Byte Magazine, 14(1):273–274, 276–278, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1989:EE

[Mee89c] B. N. Meeks. E-mail economics.

Byte Magazine, 14(4):151–155, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Meeks:1989:PPX

[Mee89d] B. N. Meeks. The protocol pack (XMODEM and its file transfer heirs evaluated). Byte Magazine, 14 (3):155–156, 158, March 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mindell:1989:DDW

[MHKO89] D. A. Mindell, J. E. Hart, S. Kittelman, and D. Ohlsen. Dealing with a digital world. Byte Magazine, 14 (8):246–248, 252, 256, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1989:LBP

[Mia89a] S. Miastkowski. Looking beyond the DOS prompt. Byte Magazine, 14(11):105–106, 108, 110–112, 114, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miastkowski:1989:NWS

[Mia89b] S. Miastkowski. A new world for DOS (DOS shell packages). Byte Magazine, 14(8):177–178, 180–181, 183, August 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mier:1984:ESE

[Mie84] Edwin E. Mier. The evolution of a standard Ethernet. Byte Magazine, 9(??):131–142, December 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miller:1984:TSS

[Mil84] Alan R. Miller. TK!Solver: Software review. Byte Magazine, 9(??): 263–272, December 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1988:OMD

[Min88] M. Minasi. OS/2’s multitasking dashboard. Byte Magazine, 13(12): 147–151, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1989:EPO

[Min89a] M. Minasi. Electing the PM (OS/2). Byte Magazine, 14(2):157– 158, February 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1989:ODG

[Min89b] M. Minasi. OS/2 disk geography.

Byte Magazine, 14(6):151–154, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1989:GYP

[Min89c] M. J. Minasi. Getting your priorities straight (OS/2). Byte Magazine, 14 (12):159–162, November 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1989:OMR

[Min89d] M. J. Minasi. OS/2 multitasking revisited. Byte Magazine, 14(13):133– 136, December 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Minasi:1989:SO

[Min89e] M. J. Minasi. The state of OS/2.

Byte Magazine, 14(11):87–88, 90, 92–93, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moshier:1986:CA

[Mos86] S. L. Moshier. Computer approximations. Byte Magazine, 11(4):161– 178, April 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McLain:1988:WOG

[MS88] M. M. McLain and T. E. Stevenson. Writing OS/2 graphics programs. Byte Magazine, 13(11):177– 184, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Modianos:1987:TIR

[MSC87] D. T. Modianos, R. C. Scott, and L. W. Cornwell. Testing intrinsic random number generators. Byte Magazine, 12(1):175–178, 1987. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McWorter:1987:CF

[MT87] McWorter and Tazelaar. Creating fractals. Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Manson:1989:RCD

[MT89a] C. Manson and K. Thurber. Remote control (distributed processing). Byte Magazine, 14(7):235–238, 240, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marshall:1989:WRP

[MT89b] T. Marshall and J. Morrill Tazelaar. Worth the RISC (personal workstations). Byte Magazine, 14(2):245– 249, February 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mundie:1989:ITI

[Mun89] C. Mundie. Interacting with the tiny and the immense (scientific visualisation). Byte Magazine, 14(4):279– 288, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Myers:1988:RWMa

[Mye88a] ?. Myers. Review of IBM Corp., IBM Personal System/2 and Personal Computer BIOS Interface Technical Reference. Byte Magazine, 13(??):??, ?? 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Myers:1988:RWMb

[Mye88b] ?. Myers. Review of Wadlow, Memory Resident Programming on the IBM PC. Byte Magazine, 13(??):??, ?? 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nadeau:1989:CGE

[Nad89] M. E. Nadeau. Cheetah Gold 33:

an economical powerhouse. Byte Magazine, 14(12):107–110, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1989:TM

[Nan89] B. Nance. TopSpeed Modula-2. Byte Magazine, 14(5):211–214, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nelson:1988:MIS

[Nel88] Theodor H. Nelson. Managing immense storage. Byte Magazine, 13 (1):224–238, January 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Newcom:1988:MCV

[New88] K. Newcom. The Micro Channel versus the AT bus. Byte Magazine, 13(11):91–98, 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nguyen:1981:NMD

[Ngu81] T. C. Nguyen. Numerical methods in data analysis. Byte Magazine, 6(??):435–446, ?? 1981. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nicholls:1988:GBP

[Nic88] B. Nicholls. Graphics: the big picture. Byte Magazine, 13(11):129– 138, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nicholls:1989:CCP

[Nic89a] B. Nicholls. The current crop (personal workstations). Byte Magazine, 14(2):235–244, February 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nicholls:1989:IRS

[Nic89b] B. Nicholls. Is it really super? (VGA standard). Byte Magazine, 14(11): 159–160, 162, 164, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nyberg:1989:IAP

[NU89] K. Nyberg and J. Udell. Integrada (ada programming support environment for the PC). Byte Magazine, 14(1):213–214, 216, 218, 220, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Obermeier:1989:TGF

[OB89] K. K. Obermeier and J. J. Barron.

Time to get fired up (neural networks). Byte Magazine, 14(8):217– 220, 222–224, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Obermeier:1987:NLP

[Obe87] K. K. Obermeier. Natural language processing. Byte Magazine, 12(??): 225–233, ?? 1987. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Obermeier:1988:SSY

[Obe88] K. K. Obermeier. Side by side; you can only simulate true parallelism on your personal computer today, but tomorrow will be another story. Byte Magazine, 13(12):275– 276, 280–283, December 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orr:1989:MES

[OGY^{+}89] K. Orr, C. Gane, E. Yourdon, P. Chen, and L. L. Constantine. Methodology: the experts speak (software engineering). Byte Magazine, 14(4):221–233, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

ONeill:1988:FTF

[O’N88] ?. O’Neill. Faster than fast Fourier. Byte Magazine, 13(??):??, ?? 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Oppenheim:1988:DDL

[Opp88a] P. Oppenheim. D the data language (DBMS). Byte Magazine, 13(10): 215–221, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Opper:1988:GT

[Opp88b] S. Opper. A groupware toolbox. Byte Magazine, 13(13):275–282, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Oren:1988:CC

[Ore88] T. Oren. The CD-ROM connection. Byte Magazine, 13(13):315–320, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pajari:1989:IAA

[Paj89] G. E. Pajari. Interrupts aren’t always best. Byte Magazine, 14(5): 261–264, May 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pascal:1989:BNW

[Pas89] F. Pascal. A brave new world? (DBMS). Byte Magazine, 14(9): 247–250, 252, 255–256, September 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pearson:1988:LDL

[Pea88] R. Pearson. Lies, damned lies, and spreadsheets (validity tests). Byte Magazine, 13(13):299–304, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perry:1986:AMA

[Per86] ?. Perry. Abstract mathematical art.

Byte Magazine, 11(??):??, ?? 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Phipps:1986:ILM

[Phi86] ?. Phipps. The inversion of large matrices; the Pan and Reif algorithm provides a solution. Byte Magazine, 11(??):??, ?? 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Phipps:1987:MPR

[Phi87] ?. Phipps. More on Pan-Reif (letter). Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Poplett:1988:DTD

[PK88] John Poplett and Rob Kurver. The DSI transputer development system. Byte Magazine, 13(2):249–254, February 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Peterson:1988:TWD

[PKAG88] R. Peterson, C. R. Killebrew, Jr., T. Albers, and K. Guttag. Taking the wraps off the 34020 (display technology). Byte Magazine, 13(9):

257–272, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1982:SST

[Pou82] J. Pournelle. Semidisk, software tools, the BDOS blues, power, and LISPs. Byte Magazine, 7(??):??, August 1982. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1984:PM

[Pou84a] D. Pountain. PROLOG on microcomputers. Byte Magazine, 9(13): 355–62, December 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1984:TSL

[Pou84b] Dick Pountain. The Transputer and its special language, Occam. Byte Magazine, 9(8):361–366, August 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pournelle:1984:R

[Pou84c] J. Pournelle. On the road. Byte Magazine, 9(10):363–382, September 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1985:CC

[Pou85a] Dick Pountain. Computers as consultants. Byte Magazine, 10(10): 367–376, October 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1985:PP

[Pou85b] Dick Pountain. Parallel processing.

Byte Magazine, 10(5):385, 386, 388, 390, 392, 395, May 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1986:PS

[Pou86a] Dick Pountain. Personal supercomputers. Byte Magazine, 11(7):363– 368, July 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1986:TM

[Pou86b] Dick Pountain. Turbocharging Mandelbrot. Byte Magazine, 11(9):359– 366, September 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1988:PP

[Pou88a] D. Pountain. Parallelizing Prolog. Byte Magazine, 13(12):387–394, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1988:ROC

[Pou88b] D. Pountain. Rekursiv: an objectoriented CPU. Byte Magazine, 13 (12):340–349, November 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1988:UPS

[Pou88c] D. Pountain. Untangling Pascal strings. Byte Magazine, 13(13):307– 314, December 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1988:PT

[Pou88d] Dick Pountain. A personal Transputer. Byte Magazine, 13(??):303– 304 (??), June 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1988:ROO

[Pou88e] Dick Pountain. Rekursiv: An object-oriented CPU. Byte Magazine, 13(??):341–349, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1989:DP

[Pou89a] D. Pountain. Digital paper. Byte Magazine, 14(2):274–280, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1989:HSM

[Pou89b] D. Pountain. HM Systems’ Minstrel workstation. Byte Magazine, 14(12):80IS–3–10, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1989:MPP

[Pou89c] D. Pountain. Minicomputer performance on a personal computer (coprocessor board). Byte Magazine, 14(8):801S/3–8, August 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1989:XWS

[Pou89d] D. Pountain. The X Window System. Byte Magazine, 14(1):353–355, 336–360, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1989:CPP

[Pou89e] Dick Pountain. Configuring parallel programs – part 1. Byte Magazine, 14(13):349–352, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1989:OI

[Pou89f] Dick Pountain. Occam II. Byte Magazine, 14(10):279–284, October 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pollack:1986:INL

[PW86] J. Pollack and D. L. Waltz. Interpretation of natural language. Byte Magazine, 11(??):189–198, February 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Quirk:1989:CNT

[Qui89a] K. Quirk. Color by numbers (thermal printer). Byte Magazine, 14(7): 177–179, July 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Quirk:1989:LLP

[Qui89b] K. Quirk. The language of lasers (printer control languages). Byte Magazine, 14(11):203–206, 208, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rasmussen:1984:QS

[Ras84] E. H. Rasmussen. Queue simulation. Byte Magazine, 9(3):157–174, March 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1988:DPT

[Ras88a] Wayne Rash, Jr. Do productivity tools help productivity? Byte Magazine, 13(12):135–137, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1988:LBW

[Ras88b] Wayne Rash, Jr. Light, bright and white (laptop screens). Byte Magazine, 13(13):321–324, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1988:OCP

[Ras88c] Wayne Rash, Jr. The odd couple (portable computers). Byte Magazine, 13(10):185–190, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1988:QW

[Ras88d] Wayne Rash, Jr. A quintet of worms. Byte Magazine, 13(??):146– 151, February 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1989:CHH

[Ras89a] Wayne Rash, Jr. Computing in hand (hand-held computers). Byte Magazine, 14(5):195–196, 198, 200, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1989:GBL

[Ras89b] Wayne Rash, Jr. Getting into bigger LANs. Byte Magazine, 14(2): 145–147, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1989:GGG

[Ras89c] Wayne Rash, Jr. Groping for groupware (group productivity software). Byte Magazine, 14(4):135– 138, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1989:JFF

[Ras89d] Wayne Rash, Jr. Just a few fax (PC fax boards). Byte Magazine, 14(6): 143–144, June 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rash:1989:MYD

[Ras89e] Wayne Rash, Jr. So, maybe you do need a LAN. Byte Magazine, 14 (1):135–136,138, January 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reed:1989:QT

[Ree89] M. Reed. The quantum transistor.

Byte Magazine, 14(5):275–281, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Reno:1989:PUO

[Ren89] E. Reno. For power users only (lotus 1-2-3 release 3.0). Byte Magazine, 14 (12):255–262, November 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robson:1981:OOS

[Rob81] D. Robson. Object-oriented software systems. Byte Magazine, 6(??): 74, August 1981. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Roberts:1987:OCH

[Rob87a] ?. Roberts. Optimizing compilers: How compilers produce fast code, and how they could be improved. Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1987:SAC

[Rob87b] P. Robinson. The sum: an AI coprocessor. Byte Magazine, 10(6):??, June 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1989:ERP

[Rob89a] P. Robinson. Easy reading (pagerecognition systems). Byte Magazine, 14(5):203–204, 206, 208, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1989:OGS

[Rob89b] P. Robinson. Opus I (graphic/ database software). Byte Magazine, 14(1):233–234, 236–237, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1989:VS

[Rob89c] P. Robinson. Variations on a screen. Byte Magazine, 14(4):251– 264, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Robinson:1989:AYP

[Rob89d] Philip Robinson. Art + 2 years = science (personal workstations). Byte Magazine, 14(2):255– 264, February 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rosenberg:1989:CDS

[Ros89] S. Rosenberg. Corel Draw shows great promise (graphics program). Byte Magazine, 14(6):213–216, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rothchild:1983:OMM

[Rot83] Edward Rothchild. Optical-memory media. Byte Magazine, 7(??):86– 106, March 1983. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rockett:1987:KA

[RS87a] A. M. Rockett and J. C. Stevenson. Karmarkar’s algorithm. Byte Magazine, 12(10):147–160, September 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rockett:1987:KAM

[RS87b] A. M. Rockett and J. C. Stevenson.

Karmarkar’s algorithm: A method for solving large linear programming problems (technical). Byte Magazine, 12(10):146–162, October 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rubel:1989:DIA

[Rub89a] M. Rubel. dBASE IV arrives. Byte Magazine, 14(2):217–222, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rubel:1989:AFN

[Rub89b] M. C. Rubel. Alpha Four: no programming required (related DBMS). Byte Magazine, 14(12):265–268, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Saal:1989:LTL

[Saa89] H. Saal. Looking for trouble (LAN performance). Byte Magazine, 14 (1):259–265, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Saffari:1989:PDW

[Saf89] B. Saffari. Putting DSPs to work. Byte Magazine, 14(13):259–272, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sartore:1989:HP

[Sar89] R. Sartore. The 80486: a hardware perspective. Byte Magazine, 14(11): 67–70, 72–74, 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Satchell:1989:DDA

[Sat89a] S. Satchell. Downsizing the desktop (ADC Powerlite 386 SX). Byte Magazine, 14(13):179–182, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Satchell:1989:XPP

[Sat89b] S. Satchell. X.25 pads performance (packet assembler/disassembler). Byte Magazine, 14(12):233–235, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1988:FFD

[SB88] G. Smarte and N. M. Baran. Face to face (display technology). Byte Magazine, 13(9):243–252, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schrodt:1984:SWV

[Sch84] P. A. Schrodt. Simulation of weighted voting: The Banzhaf index. Byte Magazine, 9(3):138–154, March 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schalkoff:1986:M

[Sch86a] Robert J. Schalkoff. MuLISP-86.

Byte Magazine, 11(??):249–252, October 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schmucker:1986:MAF

[Sch86b] K. Schmucker. MacApp: an application framework. Byte Magazine, 11 (8):189–193, August 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schrodt:1986:PIE

[Sch86c] P. A. Schrodt. Predicting international events. Byte Magazine, 11 (??):177–192, ?? 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schroeder:1986:PMS

[Sch86d] Peter B. Schroeder. Plotting the Mandelbrot set. Byte Magazine, 11 (13):207–211, December 1986. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schulman:1988:EOL

[Sch88] A. Schulman. Exploring OS/2 with a Lisp interpreter. Byte Magazine, 13(11):201–206, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Scherr:1989:PPV

[Sch89a] I. Scherr. Pepperoni and paperwork (voice/document delivery system). Byte Magazine, 14(13):309– 316, December 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schnapp:1989:CAG

[Sch89b] M. Schnapp. Clipper applications get SQL. Byte Magazine, 14(13): 211–214, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schulman:1989:DKC

[Sch89c] A. Schulman. A different kind of CASE tool. Byte Magazine, 14(5):

217–220, May 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schulman:1989:MOJ

[Sch89d] A. Schulman. Modula-2 and OS/2 join forces. Byte Magazine, 14(8): 171–174, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VonStroh:1988:LWD

[SD88] R. Von Stroh and B. Dolinar. Lighting the way (EL displays). Byte Magazine, 13(9):275–280, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1984:FLA

[Sha84] E. Shapiro. A first look at Dayflo.

Byte Magazine, 9(3):81–86, March 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1988:ACM

[Sha88a] N. C. Shammas. Ada comes to the Mac (compiler package). Byte Magazine, 13(9):213–216, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1988:BR

[Sha88b] N. C. Shammas. The BASIC revival. Byte Magazine, 13(9):295– 300, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1988:LAM

[Sha88c] N. C. Shammas. Life after DOS (multitasking operating systems). Byte Magazine, 13(11):143–150, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1988:STE

[Sha88d] N. C. Shammas. Slick (text editing environment). Byte Magazine, 13 (13):215–221, December 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1988:SCC

[Sha88e] N. C. Shammas. Smalltalk a la C (C-talk 1.0). Byte Magazine, 13(10): 201–204, October 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1988:SHD

[Sha88f] N. C. Shammas. Software for hardware-style debugging (soft-ICE debugger). Byte Magazine, 13(9): 219–222, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1988:PS

[Sha88g] E. Shapiro. Portable software. Byte Magazine, 13(12):127–132, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shammas:1989:QCM

[Sha89a] N. C. Shammas. QuickBASIC comes to the Macintosh. Byte Magazine, 14(1):223–224, 226, 228, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1989:TNW

[Sha89b] E. Shapiro. Then and now in word processing. Byte Magazine, 14(6): 137–140, June 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1989:TMP

[Sha89c] E. Shapiro. Trackball madness (Pro-Point, Works 2.0 and WorkPlus Spell on Macintosh). Byte Magazine, 14(3):129–130, 132, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1982:DSU

[SIK^{+}82] D. C. S. Smith, C. Irby, R. Kimball, B. Verplank, and E. Harlem. Designing the Star user interface. Byte Magazine, 7(4):242–282, April 1982. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1989:DRP

[SM89] B. Smith and R. Mitchell. DEC’s RISC powerhouse. Byte Magazine, 14(12):201–206, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1983:PKC

[Smi83] John Smith. Public key cryptography: An introduction to a powerful cryptographic system for use on microcomputers. Byte Magazine, 7 (??):198–218, January 1983. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). This is a simple exposition of public key cryptography.

Smith:1989:AP

[Smi89a] B. Smith. AIX on the PS/ 2s. Byte Magazine, 14(11):95–96, 98, 100– 101, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1989:HUP

[Smi89b] B. Smith. High-fashion Unix on a PC (HP display server). Byte Magazine, 14(13):205–208, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smith:1989:UC

[Smi89c] B. Smith. The Unix connection. Byte Magazine, 14(5):245–251, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Snader:1987:LFH

[Sna87] Jon C. Snader. Look it up faster with hashing. Byte Magazine, 12 (1):128–144, January 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Smarte:1989:SI

[SP89] G. Smarte and W. Penney. Sounds and images. Byte Magazine, 14(13): 243–256, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sokol:1980:TDG

[SS80] D. Sokol and J. Shepard. Threedimensional graphics for the Apple II. Byte Magazine, 5(11):148–150, November 1980. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sakamura:1989:TP

[SS89] K. Sakamura and R. Sprague. The TRON project. Byte Magazine, 14 (4):292–301, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stallings:1989:WOL

[Sta89] W. Stallings. When one LAN is not enough. Byte Magazine, 14(1):293– 298, January 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stevens:1985:REB

[Ste85] J. K. Stevens. Reverse engineering the brain. Byte Magazine, 10(??): 287–298, April 1985. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sterling:1987:MRP

[Ste87] L. Sterling. Mathematical reasoning — A Prolog program uses heuristic methods to solve equations. Byte Magazine, 12(10):177–180, October 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stein:1988:TC

[Ste88] Richard M. Stein. T800 and counting. Byte Magazine, 13(12):287–296, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stephenson:1989:PPM

[Ste89a] Peter Stephenson. Personal and private (microcomputer security). Byte Magazine, 14(6):285–288, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stevens:1989:MDP

[Ste89b] L. Stevens. Mac desktop presentation software. Byte Magazine, 14(4): 203–205, April 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stroustrup:1988:BC

[Str88] Bjarne Stroustrup. A better C? Byte Magazine, 13(??):??, ?? 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1988:NC

[TB88] Tom Thompson and Nick Baran. The NeXT computer. Byte Magazine, 13(12):158–175, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tello:1988:BMM

[Tel88] Ernest R. Tello. Between man and machine (user interfaces). Byte Magazine, 13(9):288–293, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tesler:1981:SE

[Tes81] Larry Tesler. The Smalltalk environment. Byte Magazine, 6(8):90– 147, August 1981. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). The special issue on Smalltalk has a very informative tutorial.

Thompson:1989:PP

[TH89] T. Thompson and F. Hayes. The portable and the powerful. Byte Magazine, 14(10):98–108, October 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1987:FME

[Tho87] T. Thompson. Fast math — A first look at Motorola’s 68882 math coprocessor. Byte Magazine, 12(12): 120–121, December 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thomas:1989:WOO

[Tho89a] D. Thomas. What’s in an object? (object-oriented programming). Byte Magazine, 14(3):231– 232, 234–236, 238, 240, 270–271, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:NS

[Tho89b] ?. Thompson. The next step. Byte Magazine, 14(??):??, ?? 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:AQC

[Tho89c] T. Thompson. Apple’s 32-bit Quick-Draw covers the spectrum. Byte Magazine, 14(7):99–102, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:BMP

[Tho89d] T. Thompson. Big Mac power in a small Mac box. Byte Magazine, 14(6):175–178, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:FS

[Tho89e] T. Thompson. Full-spectrum scanners. Byte Magazine, 14(4):189– 194, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:HPM

[Tho89f] T. Thompson. How to put 16 million colors to work (video-graphics boards). Byte Magazine, 14(13): 189–196, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:LAM

[Tho89g] T. Thompson. LAN aid: Mac booster modules. Byte Magazine, 14 (12):219–223, November 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1989:NSO

[Tho89h] T. Thompson. The next step (object-oriented programming simplifies and speeds software development with the NeXT computer’s NextStep). Byte Magazine, 14(3): 265–271, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thurber:1989:OHN

[Thu89] K. Thurber. OS/2 hits the networks. Byte Magazine, 14(1):285–291, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tittel:1989:MCI

[Tit89] E. Tittel. Making the connection (IBM PCs, Macintoshes and VAXes). Byte Magazine, 14(1):317– 320, 322–325, January 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Touretzky:1988:HLC

[Tou88] David S. Touretzky. How LISP has changed. Byte Magazine, 13(2):229– 234, February 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Touretzky:1989:WHH

[TP89] D. S. Touretzky and D. A. Pomerleau. What’s hidden in the hidden layers? Byte Magazine, 14(8):227– 228. 230–233, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tevanian:1989:MMF

[TS89] A. Tevanian, Jr. and B. Smith.

Mach: the model for future Unix (object-oriented operating system). Byte Magazine, 14(12):411–417, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thompson:1986:FRD

[TT86] B. Thompson and W. Thompson.

Finding rules in data. Byte Magazine, 11(??):149–158, November 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tuthill:1983:TUS

[Tut83] B. Tuthill. Typesetting on the UNIX system. Byte Magazine, 8(19):253– 265, October 1983. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tuten:1989:CCS

[Tut89] P. Tuten. Claris CAD (software package). Byte Magazine, 14(6): 209–210, June 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1988:CT

[Ude88] J. Udell. A C++ toolkit. Byte Magazine, 13(12):223–227, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1989:COP

[Ude89a] J. Udell. Clash of the objectoriented Pascals. Byte Magazine, 14(7):104–106, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1989:ETE

[Ude89b] J. Udell. Extensible text editors for programmers. Byte Magazine, 14(3):197–198, 200–202, 204, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1989:GLZ

[Ude89c] J. Udell. Greased lightning (zenith system). Byte Magazine, 14(6):114– 116, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1989:PPI

[Ude89d] J. Udell. The power of the press (Interleaf’s TEchnical Publishing Software). Byte Magazine, 14(12):271– 276, November 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1988:BSD

[Ung88a] J. Unger. Bucking the system (Dell System 310). Byte Magazine, 13 (10):179–183, October 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1988:SUW

[Ung88b] J. Unger. The Sun386i (Unix workstation). Byte Magazine, 13(13): 183–188, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1989:DHM

[Ung89a] J. Unger. Dead heat (microcomputer review). Byte Magazine, 14 (6):181–184, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1989:OME

[Ung89b] J. Unger. One man’s experience (Unix). Byte Magazine, 14(5):237– 238, 240–242, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Unger:1989:PSL

[Ung89c] J. Unger. A pair of sophisticated laptops (Zenith SupersPort 286 and Mitsubishi MP-286L). Byte Magazine, 14(2):189–194, February 1989.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Valdes:1989:EMP

[Val89a] R. Valdes. Ease into Mac programming SmethersBarnes Prototyper 2.1. Byte Magazine, 14(12): 241–244, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Valdes:1989:EST

[Val89b] R. Valdes. Extend (simulation toolkit). Byte Magazine, 14(4):197– 200, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Valdes:1989:SCM

[Val89c] R. Valdes. Smalltalk/V comes to the Mac. Byte Magazine, 14(6):201–204, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Valdes:1989:VTW

[Val89d] R. Valdes. A virtual toolkit for Windows and the Mac. Byte Magazine, 14(3):209–210, 212, 214, 216, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1988:AIW

[Van88a] M. L. Van Name. ALR improves on a winner (cache store). Byte Magazine, 13(12):205–210, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1988:KC

[Van88b] M. L. Van Name. Keeping up with the CPU. Byte Magazine, 13(11): 101–106, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:HCC

[Van89a] M. L. Van Name. High-tech computing, cafeteria style (Wells American CompuStar 286). Byte Magazine, 14 (4):179–184, April 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:PPD

[Van89b] M. L. Van Name. A portable with punch (dolch PAC). Byte Magazine, 14(1):189–192, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:PCO

[Van89c] M. L. Van Name. A PS/2 in channel only (tandy 5000 MC). Byte Magazine, 14(2):197–202, February 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:ALO

[VC89a] M. L. Van Name and B. Catchings.

Anatomy of a LAN operating system. Byte Magazine, 14(6):157–160, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:HSP

[VC89b] M. L. Van Name and B. Catchings. Handy scanners (PC hand-held scanners). Byte Magazine, 14(6):

187–191, June 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:LRO

[VC89c] M. L. Van Name and B. Catchings. The LAN road to OSI. Byte Magazine, 14(7):148–150, 152, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:MMS

[VC89d] M. L. Van Name and B. Catchings. MCA meets SX. Byte Magazine, 14 (10):181–184, October 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:PPP

[VC89e] M. L. Van Name and B. Catchings. The painlessly portable PC (NEC UltraLite). Byte Magazine, 14 (8):161–164, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VanName:1989:SDL

[VC89f] M. L. Van Name and B. Catchings. SQL: a database language sequel to dBASE. Byte Magazine, 14(11): 175–178, 180, 182, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

vanName:1989:SDD

[vNC89] M. L. van Name and B. Catchings. Serving up data (database server). Byte Magazine, 14(9):259– 260, 262, 264, September 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vose:1989:RGC

[Vos89] ?. Vose. Review of Gehani, C:

An Advanced Introduction, ANSI C Edition, Computer Science Press, 1988. Byte Magazine, 14(??):??, ?? 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

West:1989:WMD

[WA89] J. West and M. Anderson. Writing Macintosh device drivers. Byte Magazine, 14(6):96IS–53–60, June 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wadlow:1981:XAC

[Wad81] Thomas A. Wadlow. The Xerox Alto computer. Byte Magazine, 6(??):58– 68, September 1981. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Walker:1985:TBB

[Wal85] Paul Walker. The Transputer: a building block for parallel processing. Byte Magazine, 10(5):219–235, May 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waller:1988:PPP

[Wal88a] M. Waller. PC power. 1. Power protection. Byte Magazine, 13(10):270– 280, October 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waller:1988:PPB

[Wal88b] M. Waller. PC power. 2. Backup power. Byte Magazine, 13(12):353– 358, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Watson:1989:WL

[Wat89] R. Watson. Which LAN? Byte Magazine, 14(11):195–198, 200, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1987:ZKP

[Way87] ?. Wayner. Zero-knowledge proofs. Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1988:AWA

[Way88a] P. Wayner. It’s APT to write (abstract planning tool). Byte Magazine, 13(12):375–384, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1988:EFF

[Way88b] Peter Wayner. Error-free fractions. Byte Magazine, 13(??):??, ?? 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1989:VHR

[Way89a] P. Wayner. VLIW: heir to RISC? Byte Magazine, 14(8):259–262, August 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1989:SMM

[Way89b] Peter Wayner. Symbolic math on the Mac (Mathematica 1.0). Byte Magazine, 14(1):239–242, 144, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wegner:1989:LLO

[Weg89] Peter Wegner. Learning the language (object-oriented programming languages). Byte Magazine, 14(3):245–250, 252–253, 270–271, March 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weiner:1984:LRK

[Wei84] J. L. Weiner. The logical record keeper: PROLOG on the IBM. Byte Magazine, 9(9):125–31, September 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). QA76.5B98.

Wichmann:1987:BRN

[WH87] Brian A. Wichmann and ?. Hill.

Building a random-number generator: A Pascal routine for very-long-cycle random-number sequences. Byte Magazine, 12(??):??, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Waibel:1989:BBS

[WH89] A. Waibel and J. Hampshire. Building blocks for speech (neural networks). Byte Magazine, 14(8):235– 238, 240, 242, August 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Whitney:1984:GTP

[Whi84] C. A. Whitney. Generating and testing pseudorandom numbers. Byte Magazine, 9(11):128–129, November 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

White:1989:BT

[Whi89] G. White. A bus tour. Byte Magazine, 14(9):296–300, 302, September 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Williams:1983:LCS

[Wil83] Geoff Williams. The Lisa computer system. Byte Magazine, 8(2):33–50, February 1983. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Williams:1984:AMC

[Wil84] Geoff Williams. The Apple Macintosh computer. Byte Magazine, 9 (2):30–54, February 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilton:1985:PEG

[Wil85] R. Wilton. Programming the Enhanced Graphics Adapter. Byte Magazine, 10(??):209–220, ?? 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Williams:1988:KYP

[Wil88a] G. B. Williams. Keep your PC healthy (computer maintenance). Byte Magazine, 13(11):169–174, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilson:1988:FPS

[Wil88b] Pete Wilson. Floating-point survival kit. Byte Magazine, 13(3):217, March 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilson:1988:PPC

[Wil88c] Pete Wilson. Parallel processing comes to PCs (transputer boards for PC AT-compatibles and macintoshes). Byte Magazine, 13(12):213– 218, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wilton:1988:VVM

[Wil88d] R. Wilton. VGA video modes. Byte Magazine, 13(11):187–198, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Winston:1985:LR

[Win85] Patrick H. Winston. The LISP revolution. Byte Magazine, 10(4):209– 222, April 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Winograd:1988:WAG

[Win88] T. Winograd. Where the action is (groupware). Byte Magazine, 13(13):256A–258, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wirth:1984:HGM

[Wir84] N. Wirth. History and goals of Modula-2. Byte Magazine, 9(??): 145–152, August 1984. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Modula-2.

Wittman:1987:RWPb

[Wit87a] ?. Wittman. Review of: Snyder-Palmer, In Search of the Most Amazing Thing: Children, Education, and Computers, Addison-Wesley, 1986. Byte Magazine, 12 (??):??, ?? 1987. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wittman:1987:RWPa

[Wit87b] ?. Wittman. Review of: Wresch, A Practical Guide to Computer Uses in the English/Language Arts Classroom, Prentice-Hall, 1986. Byte Magazine, 12(??):??, ?? 1987. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wller:1989:MWP

[Wll89] M. Wller. Making waves (packet radio). Byte Magazine, 14(13):363– 366, December 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

West:1988:QLA

[WN88] J. West and D. Newton. A quick look at QuickCapture (frame grabber board and software). Byte Magazine, 13(13):199–204, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wong:1985:TL

[Won85] W. G. Wong. TLC-LISP. Byte Magazine, 10(10):287–292, October 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wong:1986:BLW

[Won86] William G. Wong. BYSO LISP and Waltz LISP. two implementations of LISP for the IBM personal computer and compatibles. Byte Magazine, 11 (??):293–296, July 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wong:1987:PSL

[Won87] William Wong. PC Scheme: A lexical LISP. Byte Magazine, 12 (??):223–226, March 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1988:DEG

[Woo88a] L. Wood. Data entry goes hightech (dataplex package). Byte Magazine, 13(9):231–234, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1988:DRA

[Woo88b] L. Wood. The database redefined (Agenda 1.0). Byte Magazine, 13 (13):223–229, December 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1988:PPM

[Woo88c] L. Wood. The promise of project management. Byte Magazine, 13 (12):180–192, November 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1988:SYS

[Woo88d] L. Wood. Suit yourself with Sprint (word processor). Byte Magazine, 13(10):223–228, October 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1988:TW

[Woo88e] L. Wood. Total word. Byte Magazine, 13(9):225–228, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1989:MIP

[Woo89a] L. Wood. Manage it with pictures (project management package). Byte Magazine, 14(13):223– 226, December 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1989:WWP

[Woo89b] L. Wood. A window on word processing. Byte Magazine, 14(5):221– 223, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wood:1989:SSU

[Woo89c] P. Wood. Safe and secure? (unix). Byte Magazine, 14(5):253–254, 256– 258, May 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Walker:1989:TED

[WP89] B. J. Walker and G. J. Popek. A transparent environment (distributed processing). Byte Magazine, 14(7):225–226, 228–231, 233, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wai:1989:TYP

[WSW89] G. Wai, M. L. Smith, and G. White. Take your pick (distributed processing). Byte Magazine, 14(7):215–218, 220, 222–223, July 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yager:1989:XWS

[Yag89] T. Yager. X Window System on the march. Byte Magazine, 14(10):201– 206, October 1989. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yancich:1989:UEM

[Yan89] D. M. Yancich. Using expanded memory. Byte Magazine, 14(11): 123–124, 126, 128, 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Zeidenberg:1987:MB

[Zei87] M. Zeidenberg. Modeling the brain.

Byte Magazine, 12(??):237–246, ?? 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).